### v1

# ZeroTuning: Unlocking the Initial Token’s Power to Enhance Large Language Models Without Training

Abstract
Training-free methods for enhancing large language models (LLMs) have attracted growing interest recently, with token-level attention tuning emerging as an interpretable and promising direction. However, existing methods typically rely on auxiliary mechanisms to identify important or irrelevant task-specific tokens, introducing potential bias and limiting applicability. In this work, we uncover a surprising and elegant alternative: the semantically empty initial token (e.g., <BOS> in Llama) serves as a powerful and underexplored control point for optimizing model behavior. Through theoretical analysis, we show that tuning the initial token’s attention sharpens or flattens the attention distribution over subsequent tokens, and its role as an attention sink amplifies this effect. Empirically, we find that: (1) tuning its attention improves LLM performance across tasks more effectively than tuning other task-specific tokens; (2) the effect follows a consistent trend across layers, with earlier layers having greater impact, but varies across attention heads, with different heads showing distinct preferences in how they attend to this token. Based on these findings, we propose ZeroTuning, a training-free approach that improves LLM performance by applying head-specific attention adjustments to this special token. Despite tuning only one token, ZeroTuning achieves higher average performance on text classification, multiple-choice QA, and multi-turn conversation tasks across models such as LLama, Qwen, and DeepSeek. For example, ZeroTuning improves Llama-3.1-8B by 11.71% on classification tasks, 2.64% on QA tasks, and raises its multi-turn score from 7.804 to 7.966. The method is also robust to limited resources, few-shot settings, long contexts, quantization, decoding strategies, and prompt variations. Our work sheds light on a previously overlooked control point in LLMs, offering new insights into both inference-time tuning and model interpretability.
1 Introduction
Recent research has explored the potential of lightweight and training-free methods to adapt pre-trained large language models (LLMs) for downstream applications. One promising direction is token-level attention tuning, which modifies the attention distribution at inference time without changing any parameters. Unlike fine-tuning [1, 2, 3, 4, 5, 6] or prompt engineering [7, 8, 9], which treat LLMs as a black box, token-level attention tuning introduces a more transparent way to guide the model’s focus by increasing attention to key tokens or downplaying irrelevant ones. These methods can even outperform prompt-based approaches in certain scenarios, such as open-domain question answering [10]. Examples include PASTA (manual token selection) [11], AutoPASTA (LLM-guided token identification) [10], and ACT (attention-thresholding-based sink token identification) [12]. Similar ideas have also been applied to vision-language models to mitigate hallucinations by increasing attention to image tokens [13, 14, 15].
While effective, these methods require external mechanisms to identify task-specific tokens, which may introduce potential bias (e.g., emphasizing misleading tokens) and limit applicability when the importance of tokens is unclear or attention maps are not accessible. This motivates a question: Is it possible to improve model performance by tuning the attention to a universal and task-agnostic token without relying on task-specific token identification?
In this paper, we reveal that the answer lies in a widely existing yet often overlooked token: the initial token (e.g., <BOS> in LLaMA). Though it is commonly researched and regarded as an attention sink [16, 17, 18, 19], prior tuning methods have largely ignored it, focusing instead on visible input tokens. Our theoretical analysis shows that tuning the attention of this special token can sharpen or flatten the attention distribution across subsequent task-relevant tokens, while preserving their relative importance. Its role as an attention sink further amplifies this effect (Section 2.1). Additionally, we demonstrate that modifying the initial token’s key or query states achieves similar effects when direct tuning of attention weights is infeasible.
We conducted a series of experiments to investigate the impact of tuning the initial token’s attention and found the following:
  1. 1.
    Despite lacking semantic meaning, the initial token acts as an effective control point for steering model behavior: tuning its attention consistently improves LLM performance more effectively than tuning any other token (Section 2.2).
  2. 2.
    The tuning effect propagates consistently across layers, with the shallow and middle layers contributing more than the deeper ones; jointly tuning all layers yields the best results (Section 2.3).
  3. 3.
    The impact also varies across attention heads: some heads respond positively to increased attention on this token, while others respond negatively. We categorize these as up-effective and down-effective heads, and show that selectively tuning them outperforms uniform tuning across all heads (Sections 2.4 and 2.5).
These findings motivate ZeroTuning, a simple and effective training-free method that recalibrates the initial token’s attention to improve the LLM’s performance without requiring task-specific token identification (Section  2.6). Experiments demonstrate that, despite tuning only one token, ZeroTuning achieves higher average performance on text classification (e.g., SST2, BoolQ), multiple-choice QA (e.g., MMLU, LogiQA), and multi-turn conversation (e.g., MT-Bench) benchmarks across LLMs including Llama-3.1-8B-Instruct, Llama-2-13B-Instruct, Qwen-2-7B, and Deepseek-R1-14B. For example, ZeroTuning improves Llama-3.1-8B by 11.71% on classification tasks, 2.64% on QA tasks, and raises its multi-turn score from 7.804 to 7.966. The method is also robust under various conditions, including limited resources, few-shot settings, long contexts, quantization, decoding strategies, and prompt variations.
Our work sheds light on a previously overlooked control point in LLMs, offering new insights into both inference-time tuning and model interpretability.
2 Unveiling the Power of the Initial Token
In this section, we delve into the impact of tuning the initial token’s attention and introduce our method. Unless otherwise specified, all experiments use the Llama-3.1-8B-Instruct model, with dataset details provided in Section 3.1.
2.1 Formalizing the Tuning Process
Figure 1: Effect of the scaling factor $\gamma\in[0,10]$ and initial token’s attention $a_{0}\in\{0,0.2,0.4,0.6,0.8\}$ on attention differences $|a_{i}^{\prime}-a_{j}^{\prime}|$ between non-initial tokens. Shaded regions highlight tuning regimes: excessive tuning ($\gamma\in[0,0.3]\cup[6.5,10]$) that may impair LLM performance, moderate sharpening ($\gamma\in[0.3,1]$) that amplifies token-level attention differences, and moderate smoothing ($\gamma\in[1,6.5]$) that reduces them. Decreasing $\gamma$ sharpens the distribution by amplifying attention differences, while increasing $\gamma$ smooths it by reducing disparities.
In decoder-only Transformers, the next token is generated using the final token as the query, attending to all preceding tokens (including itself) as keys. Consider a single-layer, single-head attention mechanism with an input sequence:
$\mathbf{X}=[x_{0},x_{1},\dots,x_{T-1}]\in\mathbb{R}^{d\times T}.$ (1)
At generation step $T$, the query is $x_{T-1}$, and the attention weights over the keys $x_{0}$ to $x_{T-1}$ form a probability vector:
$\boldsymbol{a}=[a_{0},a_{1},\dots,a_{T-1}],\quad a_{i}\geq 0,\quad\sum_{i=0}^{T-1}a_{i}=1,$ (2)
where $a_{0}$ denotes the attention of the initial token, while $a_{1}$ to $a_{T-1}$ correspond to subsequent tokens.
To control the influence of $x_{0}$, we scale its attention weight by a factor $\gamma>0$ and re-normalize:
$a_{0}^{\prime}=\frac{\gamma a_{0}}{D},\quad a_{i}^{\prime}=\frac{a_{i}}{D}\quad\text{for }i=1,\dots,T-1,$ (3)
where the normalization constant is
$D=\gamma a_{0}+\sum_{i=1}^{T-1}a_{i}=\gamma a_{0}+(1-a_{0})=(\gamma-1)a_{0}+1.$ (4)
This rescaling preserves the relative proportions among non-initial tokens:
$\frac{a_{i}^{\prime}}{\sum_{j=1}^{T-1}a_{j}^{\prime}}=\frac{\frac{a_{i}}{D}}{\sum_{j=1}^{T-1}\frac{a_{j}}{D}}=\frac{a_{i}}{\sum_{j=1}^{T-1}a_{j}},\quad\text{for }i\geq 1,$ (5)
but compresses or expands their differences as
$a_{i}^{\prime}-a_{j}^{\prime}=\frac{a_{i}-a_{j}}{D}=\frac{a_{i}-a_{j}}{(\gamma-1)a_{0}+1},\quad\text{for }i,j\geq 1.$ (6)
Intuitively, setting $\gamma>1$ amplifies $a_{0}$ and flattens the attention distribution over the remaining tokens. Conversely, $\gamma<1$ suppresses $a_{0}$ and sharpens distinctions among other tokens. When $\gamma=1$, the attention remains unchanged. As illustrated in Figure 1, the value of $a_{0}$ determines the strength of the tuning effect—larger $a_{0}$ values lead to more pronounced modulation. However, extreme values of $\gamma$ can be detrimental: overly small values degrade model performance, a behavior consistent with prior findings on attention sink studies [19], while excessively large values can also hurt performance by suppressing attention to downstream tokens toward near-zero values.
Then, we analyze how this tuning influences the final representations. Let the value states corresponding to each token be:
$\mathbf{V}=[\mathbf{v}_{0},\mathbf{v}_{1},\dots,\mathbf{v}_{T-1}]\in\mathbb{R}^{d_{v}\times T},$ (7)
where $\mathbf{v}_{i}\in\mathbb{R}^{d_{v}}$ is the value vector associated with token $x_{i}$. The attention output is a weighted sum of these value vectors. Applying the scaled attention weights yields:
$\mathbf{o}$ $=\sum_{i=0}^{T-1}a_{i}^{\prime}\mathbf{v}_{i}=\frac{\gamma a_{0}}{D}\mathbf{v}_{0}+\sum_{i=1}^{T-1}\frac{a_{i}}{D}\mathbf{v}_{i},$
$=\frac{1}{D}\left[\gamma a_{0}\mathbf{v}_{0}+\sum_{i=1}^{T-1}a_{i}\mathbf{v}_{i}\right].$ (8)
Previous studies have observed the “value-state drain” phenomenon, wherein the value state of the initial token tends to be much smaller than those of the other tokens on average [20, 21, 22, 18]. Therefore, the contribution of the first term $\gamma a_{0}\mathbf{v}_{0}$ to the final output is relatively small, and the value states of the subsequent tokens dominate the final representation. Tuning $\gamma$ adjusts the relative contributions of these value states, influencing the final representation.
2.2 The Unique Importance of the Initial Token
Figure 2: Impact of attention scaling factor $\gamma$ on different token positions across three tasks: (a) SST-2, (b) BoolQ, and (c) LogiQA. Modifying the initial token’s attention consistently yields significant accuracy improvements, often surpassing adjustments to other tokens.
Given the special role of the initial token in LLMs, we begin by addressing two key questions: (a) Can tuning its attention improve performance across downstream tasks? (b) Is this position more effective and influential than others? To investigate, we conduct a controlled experiment in which we uniformly scale the attention scores of a single token position across all heads and layers using a scaling factor $\gamma$. We evaluate the resulting performance on three downstream tasks: SST-2, BoolQ, and LogiQA. For comparison, we repeat the same procedure for other positions, including the second, third, middle ($\lfloor T/2\rfloor$), and final tokens. As shown in Figure 2, tuning the attention of the initial token consistently yields the largest and most stable performance gains across all tasks. Interestingly, performance varies with the direction of tuning: tasks like SST-2 benefit from up-scaling ($\gamma>1$), while BoolQ and LogiQA improve with down-scaling ($\gamma<1$).
Previous work has identified the initial token as an attention sink that helps prevent over-mixing of information during autoregressive generation [18, 19]. We extend this understanding by showing that tuning the attention of the initial token can reshape the attention distribution and improve model performance across tasks. For example (see Appendix L), in the SST-2 sentiment classification task, input sentences often contain both positive and negative cues. We find that LLMs tend to overemphasize isolated negative keywords, leading to biased predictions. By increasing attention on the initial token, we flatten the attention distribution, promoting a more balanced integration of information and fostering a more holistic understanding. In contrast, in the BoolQ task, where long passages must be inspected for a small number of critical cues, the model may miss relevant tokens due to excessively diffuse attention. Reducing initial tokens’ attention amplifies the relative differences between the remaining tokens, sharpening focus on salient input regions and improving response accuracy.
As discussed in Section 2.1, the higher the attention on the initial token, the greater the impact of such tuning. Given that attention sinks are typically concentrated at this position, the initial token is inherently more impactful when modulated, explaining its outsized influence relative to other tokens. Besides, unlike other tokens, the initial token is task-agnostic; adjusting its attention does not alter the relative importance of task-relevant tokens, thereby avoiding harmful biases such as inadvertently amplifying attention to irrelevant content.
2.3 Layer-wise Analysis of Initial Token Scaling
To better understand the impact of tuning the initial token’s attention, we examine how its effect varies when applied selectively across different layers. Following prior work on layer functionality in transformer-based models [23, 24], we divide the 32 layers of Llama-3.1-8B-Instruct into three groups: shallow layers (Layers 1–10), middle layers (Layers 11–21), and deep layers (Layers 22–31). We then perform independent attention scaling experiments for each group on six tasks: BoolQ, SST-2, SST-5, MR, LogiQA, and MathQA. Based on earlier findings, we apply a scaling range of $[0,1]$ for BoolQ and LogiQA, and $[1,2]$ for the remaining tasks.
Figure 3: Accuracy trends when scaling the initial token’s attention across different layer groups: shallow (Layers 1–10), middle (Layers 11–21), and deep (Layers 22–31). Different depths exhibit a consistent accuracy trend with varying magnitudes.
As shown in Figure 3, tuning the initial token’s attention yields consistent trends across all three layer groups (i.e., accuracy changes similarly with scaling regardless of depth), and jointly tuning all layers amplifies these benefits, often resulting in the highest accuracy. However, the magnitude of improvement varies. In most cases, tuning the shallow and middle layers leads to greater accuracy than tuning the deep layers. This aligns with prior findings that early and middle layers primarily contribute to representation learning and knowledge integration, whereas deep layers focus on task-specific reasoning over aggregated features [25, 26].
2.4 Analyzing the Role of the Initial Token Across Attention Heads
Figure 4: Accuracy of scaling the initial token’s attention in individual heads using $\gamma=1.5$ across (a) SST-2, (b) BoolQ, (c) MMLU, and (d) MathQA. Results reveal heterogeneous behavior among heads, motivating head-specific tuning strategies.
Unlike layers, which pass information sequentially, attention heads operate in parallel and contribute independently via concatenation. It remains unclear how they differ in response to the initial token. To investigate this, we increase the initial token’s attention of each head individually by applying $\gamma=1.5$, and evaluate the model’s performance on SST-2 and MMLU. For comparison, we also evaluate (i) no scaling ($\gamma=1$), (ii) uniform up-scaling ($\gamma=1.5$) across all heads, and (iii) uniform down-scaling ($\gamma=0.6$) across all heads.
As shown in Figure 4, attention heads exhibit distinct behaviors in response to initial token amplification. We categorize heads as up-effective if this modification improves performance, and down-effective if it results in performance degradation. Interestingly, the relative proportions of up-effective and down-effective heads vary across datasets, which in turn explains the observed differences in response to uniform scaling. For example, SST-2 contains more up-effective heads and thus benefits from uniform up-scaling. In contrast, MMLU has a higher proportion of down-effective heads, making down-scaling more effective than up-scaling.
These findings align with prior work suggesting that attention heads specialize into distinct functional roles during pretraining [27]. Such roles include global retrieval, structural parsing, option discrimination, and sensitivity to negation. We hypothesize that these functional differences drive the varying impact of initial token attention scaling. While some heads rely on broad attention to support global reasoning, others benefit from focused attention on salient tokens. Together, these specialized mechanisms enable language models to better understand and process diverse input patterns.
2.5 Evaluating Head-Specific Tuning Strategies
Figure 5: Accuracy comparison of different tuning strategies on (a) SST-2 and (b) MMLU. Head-specific tuning (Up, Down) consistently outperforms uniform scaling, validating the importance of accounting for head-level behavior.
Given the diversity in head responses, we investigate whether head-specific tuning offers greater effectiveness than uniform tuning. Specifically, we compare four strategies: (i) uniform scaling of all heads (All), (ii) scaling only up-effective heads (Up), (iii) scaling only down-effective heads (Down), and (iv) a hybrid strategy (Up+Down) that scales up-effective heads to a fixed optimal value and tunes down-effective ones.
As shown in Figure 5, head-specific tuning (Up, Down) yields higher accuracy and faster convergence compared to uniform scaling (All). Notably, Up is most effective when $\gamma>1$, while Down excels when $\gamma<1$. Interestingly, the Up+Down strategy does not outperform Up or Down individually, possibly due to the concatenative nature of attention heads and suboptimal joint scaling.
2.6 ZeroTuning Methodology
Building on our empirical findings, we propose ZeroTuning, a simple yet effective method that adjusts the initial token’s attention in a head-specific manner to enhance LLM performance, without requiring task-specific token identification. ZeroTuning comprises three steps:
  1. 1.
    Head Behavior Profiling: Using a small calibration set (e.g., a few validation examples), we assess each attention head’s sensitivity to scaling the initial token’s attention. A head is classified as up-effective if increased attention improves performance, and down-effective otherwise.
  2. 2.
    Selective Rescaling: Based on the dominant head type (i.e., the group with the highest proportion), we apply a scaling factor $\gamma$ exclusively to that group to adjust the initial token’s attention. [^1] We explore the effect of scaling different subsets of heads; results (Figure 8, Appendix F) show that tuning the top 20% to 70% yields better performance.
  3. 3.
    Renormalization: The scaled attention scores are re-normalized using the softmax function to maintain a valid distribution.
For optimized attention implementations (e.g., SDPA, Flash Attention) where direct modification of attention scores is infeasible, ZeroTuning applies scaling to the query or key states before attention computation. We demonstrate that tuning key states yields similar effects to attention-based tuning, but induces steeper changes in the attention distribution of subsequent tokens (see Appendix A). The following sections evaluate the performance and robustness of ZeroTuning across different tasks and model settings.
3 Experimental Results
3.1 Experimental Setup
Models, Tasks, and Datasets. Models: We evaluate ZeroTuning on four LLMs with distinct attention implementations: Llama-3.1-8B-Instruct [28] and Llama-2-13B-Chat [29] with eager attention, Qwen-2-7B [30] with SDPA attention, and DeepSeek-R1-14B [31] with Flash attention. [^2] Eager, SDPA, and Flash are official attention implementations in modern Transformer libraries. Eager computes the full attention map; SDPA uses PyTorch’s efficient API to select the optimal implementation; Flash relies on fused CUDA kernels from the FlashAttention library. Tasks and Datasets: Our experiments encompass three task types across 15 datasets: (1) Text Classification and Reasoning, including SST-2 (binary sentiment classification) [32], SST-5 (fine-grained sentiment analysis) [32], MR (movie review polarity detection) [33], SUBJ (subjectivity classification) [34], TREC (question type classification) [35], CB (commitment detection) [36], and BoolQ (boolean question answering) [37]; (2) Domain-Specific Multiple-Choice, including MMLU (cross-domain knowledge testing) [38], AQUA (math word problems) [39], MathQA (algebraic reasoning) [40], LogiQA (logical reasoning) [41], CQA (commonsense reasoning) [42], PIQA (physical commonsense QA) [43], and ARCC (scientific reasoning) [44]; and (3) Multi-Round Conversation, using MT-Bench [39].
Baselines and Evaluation Metrics.
Baselines: We benchmark ZeroTuning against three baselines: (1) vanilla inference, which performs standard inference without any modifications; (2) ACT [12], which identifies sink tokens using an attention score threshold and reduces their attention weights; and (3) Auto-PASTA [10], which leverages an LLM to locate important tokens and enhance their attention weights. For ACT, we use the original hyperparameters ($\beta=0.4$) and filter attention heads on a 10% validation set. For AutoPASTA, the evaluated LLM identifies key input regions based on prompts provided in Appendix J. Since ACT relies on explicit attention maps, it is only evaluated on LLaMA-3.1-8B-Instruct with eager attention implementation. Evaluation Metrics: We assess performance using accuracy for text classification and multiple-choice tasks. For the multi-round conversation task, we report average quality scores as evaluated by GPT-4, following the methodology outlined in [39].
Implementation Details. ZeroTuning is implemented using PyTorch and the Hugging Face Transformers library. We select zero-shot setting with standard greedy decoding for baseline consistency, and use prompt templates from [45, 46, 47] across datasets (see Appendix J for details). Unless otherwise specified, we tune the top 40% of the heads. Each experiment uses a randomly selected validation set of 500 samples with the random seed fixed to 42.
Table 1: Performance Comparison of Classification Tasks Across Models. The best performance in each dataset is bolded, the second best is underlined, and the ZeroTuning method is highlighted in gray.
Model Method Datasets SST-2 SST-2 MR BoolQ CB TREC SUBJ Avg. Llama-3.1-8B-Instruct Vanilla 73.20 45.40 89.20 69.60 82.14 14.00 44.60 59.59 ACT 85.00 43.80 90.80 58.60 82.14 15.80 44.60 60.11 Auto-PASTA 89.60 47.20 91.40 72.60 83.93 16.00 45.40 63.73 ZeroTuning 91.60 52.00 92.00 82.40 89.29 26.20 66.60 71.44 Qwen-2-7B (SDPA) Vanilla 78.80 45.40 72.40 85.00 78.50 12.60 13.00 55.10 ACT / / / / / / / / Auto-PASTA 89.00 47.00 77.70 85.00 89.29 14.00 57.00 65.57 ZeroTuning 89.60 47.20 87.40 86.40 85.71 26.60 54.40 68.19 Deepseek-R1-14B (Flash) Vanilla 91.20 49.40 89.20 83.40 89.29 20.80 50.40 67.67 ACT / / / / / / / / Auto-PASTA 92.00 52.20 89.80 83.40 92.86 22.60 50.40 69.04 ZeroTuning 93.00 51.20 90.20 88.00 92.86 32.00 55.80 71.87
Table 2: Performance Comparison of Multiple-Choice Tasks Across Models
Model Method Datasets
MMLU AQUA MathQA LogiQA CQA PIQA ARCC Avg.
Llama-3.1-8B-Instruct Vanilla 67.40 25.69 33.60 39.40 77.60 83.60 84.62 58.84
ACT 67.60 29.64 33.60 38.00 77.60 83.00 84.62 59.15
Auto-PASTA 67.00 31.23 35.20 40.40 78.20 84.60 84.62 60.18
ZeroTuning 68.80 30.43 36.60 42.80 80.40 85.40 85.95 61.48
Qwen-2-7B (SDPA) Vanilla 69.80 36.76 39.20 45.00 78.80 85.20 86.96 63.10
ACT / / / / / / / /
Auto-PASTA 69.80 39.13 39.20 45.00 82.60 85.40 86.96 64.01
ZeroTuning 70.40 39.92 40.20 47.40 81.80 86.20 87.96 64.84
Deepseek-R1-14B (Flash) Vanilla 66.60 38.74 38.20 27.80 78.20 84.20 86.62 60.05
ACT / / / / / / / /
Auto-PASTA 66.60 38.74 39.40 28.20 78.20 84.40 86.62 60.31
ZeroTuning 70.00 39.13 39.80 35.60 78.60 85.00 87.29 62.20
3.2 Overall performance of ZeroTuning
Text Classification
We first evaluate ZeroTuning on various text classification datasets using different LLMs, as shown in Table 1. Despite tuning only a single token, ZeroTuning consistently outperforms baselines and methods that require tuning more tokens. With Llama-3.1-8B-Instruct, it achieves an average improvement of 11.71% over vanilla, with peaks of 22.00% on SUBJ and 18.40% on SST-2. It outperforms AutoPASTA by an average of 7.71%. On Qwen-2-7B, ZeroTuning gains 13.09%, and on Deepseek-R1-14B, it improves by 4.20%, with a notable increase of 11.20% on TREC.
Domain-Specific Multiple Choice
Next, we evaluate ZeroTuning on common domain-specific multiple-choice datasets under various settings, as shown in Table 2. For Llama-3.1-8B-Instruct, it increases the average accuracy by 2.64%, with gains of 3.40% on LogiQA and 1.40% on MMLU. Qwen-2-7B gains 1.74%, and Deepseek-R1-14B gains 2.15%, with an outstanding 7.80% on LogiQA.
Multi-Round Conversation
We further demonstrate ZeroTuning’s effectiveness in multi-round conversations using MT-Bench [48], with results in Table 3. For Llama-3.1-8B-Instruct, ZeroTuning improves the average score by 0.162 points (7.966 vs. 7.804). For Llama-2-13B-Chat, it achieves a 0.266 points gain (6.916 vs. 6.650), showing its effectiveness in interactive settings
Table 3: MT-Bench Performance Scores for Multi-Round Conversation Across Models
Model First Turn Second Turn Average
gpt-4 8.956 9.025 8.991
Llama-3.1-8B-ZeroTuning 8.294 (+0.029) 7.638 (+0.282) 7.966 (+0.162)
Gpt-3.5-turbo 8.075 7.813 7.944
claude-instant-v1 7.800 8.013 7.906
claude-v1 8.150 7.650 7.900
Llama-3.1-8B-vanilla 8.265 7.353 7.804
Llama-2-13B-Chat-ZeroTuning 7.106 (+0.043) 6.725 (+0.487) 6.916 (+0.266)
Llama-2-13B-Chat-vanilla 7.063 6.238 6.650
4 Further Analysis
We provide additional analyses to deepen the understanding of ZeroTuning’s behavior and practical utility under varying conditions. (1) Tuning Different Matrices: Appendix A begins with a comparison between tuning attention scores and key states. Tuning attention scores yields stable improvements due to its linear influence, whereas tuning key states is sensitive and can cause sharp performance drops because of exponential effects. (2) Resource Constraints: Appendix B evaluates ZeroTuning across three simulated computational budgets, revealing consistent gains from minimal scaling (Level 0) to full head classification and parameter search (Level 2). (3) Context Length Sensitivity: Appendix C tests robustness to long contexts with irrelevant tokens, showing ZeroTuning maintains accuracy where vanilla models degrade. (4) Few-Shot Learning Robustness: Appendix D demonstrates stable few-shot gains and fewer invalid outputs, indicating ZeroTuning complements in-context learning. (5) Prompt Variation Sensitivity: Appendix G assesses prompt modifications such as dropped or altered instructions; ZeroTuning sustains strong performance. (6) Effect of Attention Head Subset Selection: Appendix F shows that tuning a moderate proportion of attention heads (20%–70%) outperforms both full and minimal tuning. (7) Quantization Robustness: Appendix H shows a consistent trend across 16-, 8-, and 4-bit quantization, with ZeroTuning partially mitigating degradation from low-precision quantization. (8) Decoding Strategy Impact: Appendix E evaluates various decoding methods (Top-$k$, Top-$p$, Beam Search), confirming ZeroTuning’s effectiveness.
5 Related Work
5.1 Token-Level Attention Tuning
Token-level attention tuning typically aims to increase attention to critical input tokens or decrease attention to less informative tokens. [49] proposes a mask perturbation method to adjust attention weights for key tokens, thereby improving translation quality. [11] introduce PASTA, which allows manual designation of important tokens during inference. This is extended by AutoPASTA [10], which uses LLMs to autonomously identify salient tokens and increase attention to them. In contrast, ACT [12] reduces attention to semantically trivial sink tokens and redirects it to meaningful content. Similar strategies have been applied to VLMs to mitigate hallucinations. PAI [13] enhances attention to image tokens at inference time to counteract text-dominant bias. IBD [14] and OPERA [15] further refine this idea by prioritizing visual information or penalizing overconfident summary tokens. While effective, these methods depend on identifying task-specific tokens, which may introduce bias (e.g., overemphasizing misleading tokens) and limit applicability when token importance is unclear or attention maps are unavailable. In contrast, our method focuses on a task-invariant initial token, removing the need for costly token identification, and can be easily applied by tuning key states.
5.2 The Magic of the Initial Token
Recent studies highlight the significance of the initial token, especially through the lens of the attention sink phenomenon, where it draws substantial attention despite low semantic content. [16] show that preserving such tokens is critical for maintaining performance in sliding window attention. [17] attribute this effect to softmax normalization and causal masking, while [18] and [19] identify architectural biases that amplify attention to the initial token, including key-query alignment and LayerNorm effects. Functionally, the initial token is hypothesized to serve as a stabilizing “no-op” anchor, enhancing robustness to prompt variations [19]. It has been leveraged in applications such as long-context modeling [50, 16], but also poses challenges for quantization due to its high attention weight [2, 51]. While previous work has identified the structural and functional importance of the initial token, its potential as a target for attention tuning remains underexplored. In this work, we provide a detailed analysis of attention tuning of the initial token across layers and heads, demonstrating its consistent influence across different tasks. Our approach bridges the gap between the these lines of research by proposing a novel method that advances interpretable attention tuning.
6 Conclusion
In this work, we propose ZeroTuning, a novel training-free method to enhance large language models (LLMs) by adjusting the attention of the initial token. Our findings reveal that tuning the initial token’s attention can lead to substantial improvements in LLM performance across a variety of tasks, including text classification, question answering, and multi-round conversations. Notably, ZeroTuning outperforms previous task-specific tuning methods, achieving robust and transferable gains without requiring task-specific token identification. The approach offers a fresh perspective on optimizing LLMs by leveraging the often-overlooked power of the initial token, making it a simple yet effective solution for enhancing model performance. This work advances inference-time tuning and contributes to the interpretability of LLMs. It opens up new avenues for further exploration and optimization in the field.
Appendix A The Effect of Tuning Different Matrices
Figure 6: Accuracy of tuning the initial token’s attention scores and key states over (a) SST-2, (b) BoolQ, (c) AQUA, and (d) LogiQA.
In certain scenarios where the attention map is not explicitly computed, it is challenging to influence the final representation by modifying the attention weights. Therefore, we consider tuning the key or query states as an alternative approach. As illustrated in the Figure 6, we observe that within an appropriate scaling range, tuning the key state exhibits a similar trend to tuning the attention score. However, we find that directly tuning the key states is more sensitive: when the scaling factor is too small or too large, the performance of the LLM drops sharply, while tuning the attention score results in more stable performance.
We now analyze the theoretical differences between applying the scaling factor $\gamma$ to the attention scores versus the key states. To begin, we revisit and extend the attention weight formulation from Section 2.1. For a sequence of length $T$, the attention weight for token $i$ is given by:
$a_{i}=\frac{\exp(z_{i})}{\sum_{m=0}^{T-1}\exp(z_{m})},$ (9)
where $z_{i}$ denotes the logit for token $i$, given by:
$z_{i}=\frac{\mathbf{q}^{\top}\mathbf{k}_{i}}{\sqrt{d_{k}}},$ (10)
with $\mathbf{q}\in\mathbb{R}^{d_{k}}$ as the query vector, $\mathbf{k}_{i}\in\mathbb{R}^{d_{k}}$ as the key vector for token $i$, and $d_{k}$ as the dimensionality of the key vectors. Note that $a_{0}$ corresponds to the initial token, and $\sum_{i=0}^{T-1}a_{i}=1$.
Tuning the Attention Score
As derived in Section 2.1, when tuning the attention score, the difference between the attention weights of non-initial tokens $i,j\geq 1$ becomes:
$a_{i}^{\prime}-a_{j}^{\prime}=\frac{a_{i}-a_{j}}{D}=\frac{a_{i}-a_{j}}{(\gamma-1)a_{0}+1}.$ (11)
Next, we expand $a_{0}$, $a_{i}$, and $a_{j}$ as follows:
$a_{i}^{\prime}-a_{j}^{\prime}$ $=\frac{a_{i}-a_{j}}{(\gamma-1)a_{0}+1}$
$=\frac{\exp(z_{i})-\exp(z_{j})}{\left(\sum_{k=0}^{T-1}\exp(z_{k})\right)\left[(\gamma-1)\frac{\exp(z_{0})}{\sum_{k=0}^{T-1}\exp(z_{k})}+1\right]}$
$=\frac{\exp(z_{i})-\exp(z_{j})}{(\gamma-1)\exp(z_{0})+\sum_{k=0}^{T-1}\exp(z_{k})}.$ (12)
Tuning the Key State
Now, consider scaling the key state by $\gamma$, i.e., $\mathbf{k}_{0}^{\prime}=\gamma\mathbf{k}_{0}$. This changes the logit for the initial token:
$z_{0}^{\prime}=\frac{\mathbf{q}^{\top}(\gamma\mathbf{k}_{0})}{\sqrt{d_{k}}}=\gamma z_{0},$ (13)
while the logits for other tokens remain unchanged: $z_{i}^{\prime}=z_{i}$ for $i\geq 1$. The tuned attention weights are then:
$a_{i}^{\prime}=\frac{\exp(z_{i}^{\prime})}{\sum_{j=0}^{T-1}\exp(z_{j}^{\prime})}=\frac{\exp(z_{i})}{\exp(\gamma z_{0})+\sum_{j=1}^{T-1}\exp(z_{j})},\quad\text{for }i\geq 1.$ (14)
The attention difference for non-initial tokens $i,j\geq 1$ is derived as:
$a_{i}^{\prime}-a_{j}^{\prime}$ $=\frac{\exp(z_{i})}{\exp(\gamma z_{0})+\sum_{k=1}^{T-1}\exp(z_{k})}-\frac{\exp(z_{j})}{\exp(\gamma z_{0})+\sum_{k=1}^{T-1}\exp(z_{k})}$
$=\frac{\exp(z_{i})-\exp(z_{j})}{\exp(\gamma z_{0})+\sum_{k=1}^{T-1}\exp(z_{k})}.$ (15)
The denominator in equation 12 includes the linear term $(\gamma-1)\exp(z_{0})$ of $\gamma$, whereas the denominator in equation 15 contains the exponential component $\exp(\gamma z_{0})$. This indicates that tuning the attention weights results in a smoother effect, while tuning the key states has a more abrupt impact.
Appendix B Performance Under Resource Constraints
Computational constraints are common in real-world applications and can limit the feasibility of head classification and parameter optimization in LLMs. To investigate how well ZeroTuning adapts to such conditions, we define three resource constraint levels based on available computational resources:
  • Level 0: Severely limited resources that prevent both head classification and parameter search.
  • Level 1: Moderately limited resources that allow parameter search but not head classification.
  • Level 2: Ample resources that support both head classification and parameter search.
We evaluate ZeroTuning’s performance across these levels using the LLaMA-3.1-8B model. At Level 0, we apply fixed scaling factors ($\gamma=2$ and $\gamma=0.6$) to all attention heads, reflecting dataset-specific scaling preferences as explored in Section 2.2. Additionally, we introduce a hybrid approach at Level 0, which selects the best-performing $\gamma$ for each dataset. At Level 1, ZeroTuning uses uniform scaling across all heads with an optimized $\gamma$. At Level 2, it classifies attention heads, scales only the over-mixing or under-mixing heads, and searches for the optimal $\gamma$.
Table 4: Performance of ZeroTuning Under Different Resource Constraints
Method Classification Avg. Class. Multiple Choice Avg. MC
SST-2 SST-5 MR BoolQ CB TREC SUBJ MMLU AQUA MathQA LogiQA CQA PIQA ARCC
Vanilla 73.20 45.40 89.20 69.60 82.14 14.00 44.60 59.73 67.40 25.69 33.60 39.40 77.60 83.60 84.62 58.84
Level 0 ($\gamma=2$) 86.20 49.20 91.00 70.06 82.41 12.00 44.80 62.24 65.60 28.46 34.40 37.40 78.20 82.40 83.61 58.58
Level 0 ($\gamma=0.6$) 53.80 43.40 82.40 72.00 83.93 17.20 44.60 56.76 67.00 22.53 32.60 40.20 77.60 82.60 83.61 58.02
Level 0 (Hybrid) 86.20 49.20 91.00 72.00 83.93 17.20 44.80 63.36 67.00 28.46 34.40 40.20 78.20 82.60 83.61 59.21
Level 1 89.60 49.00 91.40 71.20 83.93 21.80 45.20 64.59 68.00 30.04 35.00 42.80 79.20 83.80 84.62 60.49
Level 2 91.60 52.00 92.00 82.40 89.29 26.20 66.60 71.44 68.80 30.43 36.60 42.80 80.40 85.40 85.95 61.48
As shown in Table 4, ZeroTuning consistently improves performance across all resource levels. Even at Level 0, where resources are tightly constrained, the hybrid approach delivers steady gains over vanilla inference. These improvements become more substantial at Levels 1 and 2, where additional resources enable parameter optimization and head classification. Specifically, compared to the vanilla baseline, ZeroTuning increases average accuracy on text classification tasks by 3.63 percentage points at Level 0 (Hybrid), 4.86 percentage points at Level 1, and 11.71 percentage points at Level 2. For multiple-choice tasks, the corresponding gains are 0.37, 1.65, and 2.64 percentage points, respectively.
Appendix C Sensitivity to Different Context Lengths
To investigate how the distance between the initial token and task-relevant tokens affects model behavior, we evaluate the sensitivity of ZeroTuning under varying context lengths. Specifically, we insert task-irrelevant tokens between the initial token and the original input to artificially extend the context. This allows us to isolate the impact of increased token distance on attention and performance.
We conduct experiments using Llama-3.1-8B-Instruct and apply ZeroTuning with the same set of heads and scaling factors used in the previous base (non-extended) context setting. This ensures that any performance change is due solely to increased context length rather than re-optimized tuning parameters.
As shown in Table 5, the performance of vanilla LLMs consistently degrades as context length increases, likely due to disrupted information mixing caused by the inserted tokens. In contrast, ZeroTuning remains robust across all tested context lengths, yielding consistent and often significant improvements even under extreme cases of context extension. These results suggest that tuning the initial token’s attention can effectively stabilize information flow, even when relevant content is pushed further away in the input sequence.
Table 5: Impact of Context Length on ZeroTuning Performance
Dataset Method Extra Context Length Average
0 100 200 300
SST-2 Vanilla 73.20 68.40 59.20 32.00 58.20
ZeroTuning 91.60 89.20 87.40 85.40 88.40
Diff 18.40 20.80 28.20 53.40 30.20
BoolQ Vanilla 69.60 68.60 67.60 68.60 68.60
ZeroTuning 82.40 81.80 81.40 81.20 81.70
Diff 12.80 13.20 13.80 12.60 13.10
LogiQA Vanilla 39.40 36.60 36.20 35.80 37.00
ZeroTuning 42.40 43.00 41.00 41.00 41.85
Diff 3.00 6.40 4.80 5.20 4.85
PIQA Vanilla 83.60 82.20 81.20 80.60 81.90
ZeroTuning 85.40 83.80 83.20 82.80 83.80
Diff 1.80 1.60 2.00 2.20 1.90
Appendix D Robustness Across Few-Shot Scenarios
Few-shot learning has become a widely adopted approach to improve the performance of LLMs by providing a small number of in-context examples, enabling adaptation to specific tasks with minimal data [52]. Building on previous zero-shot evaluations, we now evaluate the robustness of ZeroTuning in 1-shot and 2-shot scenarios across four datasets: SST-5, BoolQ, MMLU, and AQUA. To ensure consistency, we fix the randomly selected examples, maintain the selected head and scaling factor throughout the experiments.
The results in Table 6 show that ZeroTuning consistently outperforms the vanilla baseline across both 1-shot and 2-shot settings. In the 1-shot scenario, ZeroTuning achieves an average accuracy improvement of 1.85% over the vanilla model, with notable gains of 2.0% on BoolQ (82.40% vs. 80.40%) and 1.8% on SST-5 (49.40% vs. 47.60%). In the 2-shot scenario, the average improvement increases to 3.08%, with a significant 7.12% increase on AQUA (32.81% vs. 25.69%) and 2.0% on SST-5 (52.40% vs. 50.40%). Notably, ZeroTuning in the zero-shot setting outperforms vanilla few-shot baselines, achieving higher accuracy without the additional context overhead and contextual biases introduced by in-context examples.
Our results also highlight the following key findings:
  1. 1.
    ZeroTuning improves LLM performance, even when few-shot learning does not benefit the base model. Most datasets show improvements with few-shot learning, likely due to clearer patterns and better output formatting. However, some datasets, like MMLU, experience performance drops, possibly due to increased confusion from the examples. Despite this, ZeroTuning still leads to consistent performance gains.
  2. 2.
    Similar to Few-Shot Learning, ZeroTuning reduces invalid responses from LLMs, indicating improved instruction following. For instance, in the SST-2 dataset, LLMs sometimes output incorrect responses like “neutral” in zero-shot settings when they should respond with “positive” or “negative”. Few-shot learning helps the model understand the expected format, improving accuracy. Interestingly, ZeroTuning also reduces these errors, suggesting that it helps the model better understand task-relevant information.
Table 6: Comparison of Vanilla and ZeroTuning Performance Across Few-Shot Learning Scenarios
Shot Method SST-5 BoolQ MMLU AQUA Average
0-Shot Vanilla 45.4 69.6 67.4 25.7 52.0
ZeroTuning 52.0 82.4 68.80 30.4 58.40
Diff 6.6 12.8 1.4 4.7 6.4
1-Shot Vanilla 47.6 80.4 61.8 28.1 54.5
ZeroTuning 49.4 82.4 63.4 30.0 56.3
Diff 1.8 2.0 1.6 1.9 1.8
2-Shot Vanilla 50.4 83.4 64.4 25.7 56.0
ZeroTuning 52.4 85.0 66.0 32.8 59.1
Diff 2.0 1.6 1.6 7.1 3.1
Appendix E Impact of Decoding Strategies
Decoding strategies play a crucial role in shaping the output behavior of LLMs, and can influence performance across tasks. We evaluate the robustness of ZeroTuning over three strategies: Top-k Sampling, Top-p Sampling, and Beam Search, using Llama-3.1-8B on MMLU and SST-2, with results in Table 7.
Table 7: Performance Comparison Across Decoding Strategies with and without ZeroTuning on MMLU and SST-2.
Dataset Method Top-$k$ Sampling Top-$p$ Sampling Beam Search Average
MMLU Vanilla 63.80 63.80 63.00 63.53
ZeroTuning 65.80 66.00 65.20 65.67
Diff 2.00 2.20 2.20 2.13
SST-2 Vanilla 64.40 66.60 66.60 65.87
ZeroTuning 89.20 89.60 89.20 89.33
Diff 24.80 23.00 22.60 23.47
Across all decoding strategies, ZeroTuning consistently improves over the vanilla baseline. On MMLU, it yields performance gains of 2.0% with Top-$k$, 2.2% with Top-$p$, and 2.2% with Beam Search, resulting in an average improvement of 2.1%. On SST-2, the improvements are even more substantial: 24.8% with Top-$k$, 23.0% with Top-$p$, and 22.6% with Beam Search, with an average gain of 23.5%.
Appendix F The Effect of Different Numbers of Heads
As shown in Figure 8, we observe that tuning an appropriate proportion of attention heads leads to the best performance. Specifically, Figure 8a presents results on the SST-2 dataset, where we tune the up-effective heads, while Figure 8b reports performance on the MMLU dataset with the down-effective heads. Across both datasets, we find that tuning a moderate proportion of heads (typically between 40% and 70%) achieves the highest accuracy. In contrast, tuning too few or too many heads tends to degrade performance, suggesting that selective head tuning is key to effective inference-time adaptation.
Figure 7: Accuracy of tuning different proportions of heads. (a) SST-2: tuning up-effective heads; (b) MMLU: tuning down-effective heads.
Appendix G Sensitivity to Prompt Variations
Prompts play a crucial role in guiding LLM behavior and typically consist of three components: Instruction1 (task guidance), Question (the actual query), and Instruction2 (output format specification). To evaluate the robustness of ZeroTuning under prompt perturbations, we perform experiments on the LLaMA-3.1-8B model using MMLU and SST-2 under three prompt formats: Full Prompt (Instruction1 + Question + Choices + Instruction2), Drop Instruction1, and Modify Instruction2. Detailed prompt examples are provided in Appendix J.
As shown in Table 8, ZeroTuning consistently improves performance across all prompt configurations, and maintains strong performance even when key instructions are modified or omitted, demonstrating its distinctive ability to regulate and adapt to prompt variations. On MMLU, the performance gains range from 1.2% to 2.4%, with an average improvement of 1.7%. On SST-2, the gains are more substantial, ranging from 24.8% to 26.2%, with an average improvement of 25.3%.
Table 8: Effect of Prompt Variations on Performance with and without ZeroTuning.
Prompt Format Method MMLU SST-2 Average
Full Prompt Vanilla 67.40 64.40 65.90
ZeroTuning 68.80 89.20 79.00
Diff 1.4 24.8 13.1
Drop Instruction1 Vanilla 66.80 64.40 65.60
ZeroTuning 68.00 89.20 78.60
Diff 1.2 24.8 13.0
Modify Instruction2 Vanilla 61.80 61.80 61.80
ZeroTuning 64.20 88.00 76.10
Diff 2.4 26.2 14.3
Appendix H The Effect of Different Quantization Configurations
As shown in Figure 8, we observe:
(a) Quantizing to 8-bit results in only a slight accuracy decrease compared to 16-bit, while 4-bit quantization leads to a significant accuracy decrease. However, by appropriately tuning attention to the initial token, we find that the best accuracy with 8-bit quantization becomes comparable to that of the 16-bit model on the SST-2 and BoolQ datasets. This suggests that our method can partially compensate for the performance loss caused by quantization.
(c) The accuracy trends across different quantization levels are largely similar. This consistency may offer useful insights for future work, for instance, searching for optimal parameters using low-precision models and transferring them to higher-precision models.
Figure 8: Accuracy when tuning under different quantization configurations.
Appendix I Limitations
Our current approach still relies on access to the ground-truth signals and accuracy to guide the tuning direction. As a result, it cannot yet operate in a fully unsupervised manner based solely on the properties of the input or the model’s attention distribution. Although we explored the effectiveness of the method under resource-constrained scenarios, such as using fixed scaling factors or skipping exhaustive traversal and classification of attention heads, these simplified variants consistently underperformed compared to the full method.
Additionally, our experiments follow standard evaluation protocols used in prior work [45, 46, 47, 12], and due to computational and inference resource constraints, we do not evaluate performance on longer-context generation tasks involving complex reasoning or multi-step thinking.
Future work could investigate the differences in attention distributions across layers and heads more deeply, which may lead to improved unsupervised tuning strategies. In addition, it may be fruitful to explore other tuning schemes, such as dynamically adjusting attention at critical positions in the long generated sequence.
Appendix J Prompts Used for Each Dataset
Here, we list all the prompts we used in this paper on different datasets:
For multiple choice task, we use the following prompt:
Prompt for Multiple-Choice Tasks Generate the correct answer to the following question. $<$Question$>$ $<$choice 1$>$ $<$choice 2$>$ $<$choice 3$>$ Answer:"
For text classification, we use different prompts for different datasets.
Prompt for SST-2 "Classify the sentiment of the user’s message into one of the following categories: ’positive’ or ’negative’. Sentence: $<$sentence$>$ Sentiment: "
Prompt for SST-5 "Classify the sentiment of the user’s message into one of the following categories: ’terrible’, ’negative’, ’neutral’, ’positive’, or ’great’. Sentence: $<$sentence$>$ Sentiment: "
Prompt for MR "Classify the sentiment of the movie’s review into one of the following categories: ’positive’ or ’negative’. Review: $<$sentence$>$ Sentiment: "
Prompt for TREC "Classify the given questions into the following categories: ’Description’, ’Entity’, ’Expression’, ’Person’, ’Number’, or ’Location’. Question: $<$sentence$>$ Type: "
Prompt for CB "Read the following paragraph and determine if the hypothesis is true. Premise: $<$premise$>$ Hypothesis: $<$hypothesis$>$. Answer: "
Prompt for BoolQ "Read the text and answer the question by True or False. Text: $<$passage$>$ Question: $<$question$>$? Answer: "
Prompt for SUBJ "Classify the input into one of the following categories: subjective or objective. Input: $<$text$>$ Category: "
Appendix K Prompt for key tokens identification
Prompt for key tokens identification Below is a question. Please extract the key content words or phrases from the question that are crucial for understanding and answering it correctly. These are typically the nouns, verbs, adjectives, or multi-word expressions that define the subject, action, or relation in the question. Output your selection as a Python list, where each element is a word or a phrase enclosed in quotes. For example, for the question ’What is the boiling point of water?’, the key words might be [’boiling point’, ’water’]. Question: {question} Key Words:
Appendix L LLM Output Example
L.1 Examples of SST-2 Dataset
Example 1:
Classify the sentence into one of the following sentiments: positive or negative.
Sentence: “hardly a masterpiece, but it introduces viewers to a good charitable enterprise and some interesting real people.”
Sentiment:
LLM Output: negative
Ground Truth: positive
Example 2:
Classify the sentence into one of the following sentiments: positive or negative.
Sentence: “generally, clockstoppers will fulfill your wildest fantasies about being a different kind of time traveler, while happily killing 94 minutes.”
Sentiment:
LLM Output: negative
Ground Truth: positive
Example 3:
Classify the sentence into one of the following sentiments: positive or negative.
Sentence: “whether you like rap music or loathe it, you can’t deny either the tragic loss of two young men in the prime of their talent or the power of this movie.”
Sentiment:
LLM Output: neutral
Ground Truth: positive
Example 4:
Classify the sentence into one of the following sentiments: positive or negative.
Sentence: “generally, clockstoppers will fulfill your wildest fantasies about being a different kind of time traveler, while happily killing 94 minutes.”
Answer:
LLM Output: The
Ground Truth: positive
L.2 Examples of BoolQ Dataset
Example 1:
Read the text and answer the question by True or False.
Text: Countdown (game show) – The contestant in control chooses six of 24 shuffled face-down number tiles, arranged into two groups: 20 “small numbers” (two each of 1 through 10), and four “large numbers” of 25, 50, 75 and 100. Some special episodes replace the large numbers with 12, 37, 62 and 87. The contestant decides how many large numbers are to be used, from none to all four, after which the six tiles are randomly drawn and placed on the board. A random three-digit target number is then generated by an electronic machine, affectionately known as “CECIL” (which stands for Countdown’s Electronic Calculator In Leeds). The contestants have 30 seconds to work out a sequence of calculations with the numbers whose final result is as close to the target number as possible. They may use only the four basic operations of addition, subtraction, multiplication and division, and do not have to use all six numbers. A number may not be used more times than it appears on the board. Fractions are not allowed, and only positive integers may be obtained as a result at any stage of the calculation. As in the letters rounds, any contestant who does not write down their calculations in time must go first, and both contestants must show their work to each other if their results and calculations are identical.
Question: do you have to use all the numbers on countdown?
Answer:
LLM Output: True
Ground Truth: False
Example 2:
Read the text and answer the question by True or False.
Text: Hawaii Five-0 (2010 TV series, season 8) – The eighth season of the CBS police procedural drama series Hawaii Five-0 premiered on September 29, 2017 for the 2017–18 television season. CBS renewed the series for a 23 episode eighth season on March 23, 2017. However, on November 6, 2017 CBS ordered an additional episode for the season and did the same again on February 8, 2018 bringing the count to 25 episodes. The season concluded on May 18, 2018. The eighth season ranked #18 for the 2017-18 television season and had an average of 11 million viewers. The series was also renewed for a ninth season.
Question: will hawaii five o have a season 8?
Answer:
LLM Output: False
Ground Truth: True

### v2

@@
 Abstract
-Training-free methods for enhancing large language models (LLMs) have attracted growing interest recently, with token-level attention tuning emerging as an interpretable and promising direction.
-However, existing methods typically rely on auxiliary mechanisms to identify important or irrelevant task-specific tokens, introducing potential bias and limiting applicability.
-In this work, we uncover a surprising and elegant alternative: the semantically empty initial token (e.g., <BOS> in Llama) serves as a powerful and underexplored control point for optimizing model behavior.
-Through theoretical analysis, we show that tuning the initial token’s attention sharpens or flattens the attention distribution over subsequent tokens, and its role as an attention sink amplifies this effect.
-Empirically, we find that: (1) tuning its attention improves LLM performance across tasks more effectively than tuning other task-specific tokens; (2) the effect follows a consistent trend across layers, with earlier layers having greater impact, but varies across attention heads, with different heads showing distinct preferences in how they attend to this token.
-Based on these findings, we propose ZeroTuning, a training-free approach that improves LLM performance by applying head-specific attention adjustments to this special token.
-Despite tuning only one token, ZeroTuning achieves higher average performance on text classification, multiple-choice QA, and multi-turn conversation tasks across models such as LLama, Qwen, and DeepSeek.
-For example, ZeroTuning improves Llama-3.1-8B by 11.71% on classification tasks, 2.64% on QA tasks, and raises its multi-turn score from 7.804 to 7.966.
-The method is also robust to limited resources, few-shot settings, long contexts, quantization, decoding strategies, and prompt variations.
-Our work sheds light on a previously overlooked control point in LLMs, offering new insights into both inference-time tuning and model interpretability.
+Token-level attention tuning—a class of training-free methods including Post-hoc Attention Steering (PASTA) and Attention Calibration (ACT)—has emerged as a promising way to improve frozen LLMs with interpretable interventions.
+However, these methods depend on auxiliary heuristics to identify “important” task-specific tokens, which can introduce bias and limit applicability when token importance is unclear or when using optimized kernels where attention maps are inaccessible.
+We propose a simpler and more elegant alternative: acting only on the initial token (e.g., <BOS> in LLaMA).
+We show theoretically that adding lightweight biases to this token’s attention logits monotonically controls the entropy of the downstream attention distribution—an effect amplified by its natural function as an attention sink.
+Our empirical analysis reveals that this tuning process can positively affect LLMs and better unlock their pretrained knowledge, with stronger effects in early layers and distinct scaling preferences across attention heads.
+Building on these insights, we introduce ZeroTuning: a training-free method that improves LLM performance by applying head-specific attention adjustments to the initial token, requiring zero parameter updates.
+We present two variants: a supervised mode that calibrates on validation examples, and a novel unsupervised mode that directly minimizes the model’s output entropy.
+Our method requires no KV‑cache or decoding changes, and is kernel‑agnostic (works with SDPA and FlashAttention).
+The method is lightweight and requires only four lines of modification to standard LlamaAttention code.
+It achieves broad gains across 15 datasets and outperforms previous, more complex methods; for instance, with Llama-3.1-8B, it yields relative improvements of 19.9% on classification, 4.5% on question answering, and 2.1% on dialogue.
+ZeroTuning also works out-of-the-box with quantized inference and maintains its performance improvements with increasing context lengths.
+Our code and runnable demo are available at https://anonymous.4open.science/r/ZeroTuning.
 1 Introduction
-Recent research has explored the potential of lightweight and training-free methods to adapt pre-trained large language models (LLMs) for downstream applications.
-One promising direction is token-level attention tuning, which modifies the attention distribution at inference time without changing any parameters.
-Unlike fine-tuning [1, 2, 3, 4, 5, 6] or prompt engineering [7, 8, 9], which treat LLMs as a black box, token-level attention tuning introduces a more transparent way to guide the model’s focus by increasing attention to key tokens or downplaying irrelevant ones.
-These methods can even outperform prompt-based approaches in certain scenarios, such as open-domain question answering [10].
-Examples include PASTA (manual token selection) [11], AutoPASTA (LLM-guided token identification) [10], and ACT (attention-thresholding-based sink token identification) [12].
-Similar ideas have also been applied to vision-language models to mitigate hallucinations by increasing attention to image tokens [13, 14, 15].
-While effective, these methods require external mechanisms to identify task-specific tokens, which may introduce potential bias (e.g., emphasizing misleading tokens) and limit applicability when the importance of tokens is unclear or attention maps are not accessible.
-This motivates a question: Is it possible to improve model performance by tuning the attention to a universal and task-agnostic token without relying on task-specific token identification?
-In this paper, we reveal that the answer lies in a widely existing yet often overlooked token: the initial token (e.g., <BOS> in LLaMA).
-Though it is commonly researched and regarded as an attention sink [16, 17, 18, 19], prior tuning methods have largely ignored it, focusing instead on visible input tokens.
-Our theoretical analysis shows that tuning the attention of this special token can sharpen or flatten the attention distribution across subsequent task-relevant tokens, while preserving their relative importance.
-Its role as an attention sink further amplifies this effect (Section 2.1).
-Additionally, we demonstrate that modifying the initial token’s key or query states achieves similar effects when direct tuning of attention weights is infeasible.
-We conducted a series of experiments to investigate the impact of tuning the initial token’s attention and found the following:
+Training-free methods have been widely explored to enhance Large Language Models (LLMs) performance at inference time.
+Among these, token-level attention tuning has emerged as a particularly promising direction, offering an interpretable way to steer model behavior by modifying the attention distribution without any parameter updates.
+Unlike fine-tuning (Hu et al., 2021; Dettmers et al., 2023) or prompt engineering (Wei et al., 2022; Wang et al., 2022; Han et al., 2025a; b), which largely treat LLMs as black boxes, attention tuning provides a transparent mechanism to guide the model’s focus.
+Methods such as Post-hoc Attention Steering (PASTA (Zhang et al., 2023a), AutoPASTA (Zhang et al., 2024a)), and Attention Calibration (ACT) (Yu et al., 2024) have demonstrated the power of this approach, even outperforming prompting-based techniques in complex tasks like open-domain question answering (Zhang et al., 2024a).
+Similar principles have been successfully applied to vision-language models to mitigate hallucinations by re-weighting attention towards image tokens (Liu et al., 2024b; Zhu et al., 2024; Wei & Zhang, 2024).
+However, the efficacy of these methods is fundamentally constrained by their reliance on external, often heuristic, mechanisms to identify task-specific ”important” tokens.
+This dependency not only introduces the risk of bias (e.g., amplifying misleading cues) but also limits their applicability in scenarios where token importance is ambiguous or when using optimized attention kernels where attention maps are inaccessible.
+This critical limitation motivates a fundamental question: Is it possible to enhance model performance by tuning a universal, task-agnostic token, thereby bypassing the need for fragile, task-specific token identification?
+Figure 1: Overview of the ZeroTuning method and its effectiveness.
+Left: Our method requires only a few lines of code to scale the initial token’s attention within the model’s forward pass.
+Right: This simple intervention yields significant and consistent improvements across a variety of LLMs and tasks.
+In this paper, we demonstrate that the answer is a definitive yes.
+The solution lies not in adding complexity, but in leveraging a ubiquitous yet often overlooked architectural artifact: the initial token (e.g., <BOS> in LLaMA).
+While its tendency to function as an ”attention sink” is well-documented (Xiao et al., 2023; Kaul et al., 2024; Gu et al., 2024; Barbero et al., 2025), its potential as an active control lever for performance enhancement has remained largely untapped.
+Our theoretical analysis reveals that modulating the attention on this single token allows for precise control – sharpening or smoothing – over the entire downstream attention distribution while preserving the relative importance of task-relevant tokens.
+Its natural role as an attention sink further amplifies this effect (Sec 3.1).
+Furthermore, we show this control can be achieved even without direct access to attention weights by modifying the initial token’s key or query states.
+To validate this insight, we conducted a systematic investigation and uncovered three key findings:
   1. 1.
-    Despite lacking semantic meaning, the initial token acts as an effective control point for steering model behavior: tuning its attention consistently improves LLM performance more effectively than tuning any other token (Section 2.2).
+    Despite lacking semantic meaning, the initial token acts as a positive and effective control point for steering model behavior.
+Tuning its attention corrects the model’s biases and unlocks its pretrained knowledge with reduced output entropy.
+This approach consistently yields greater performance gains than tuning any other token (Sec 3.2).
   2. 2.
-    The tuning effect propagates consistently across layers, with the shallow and middle layers contributing more than the deeper ones; jointly tuning all layers yields the best results (Section 2.3).
+    The influence of this tuning is systemic, propagating consistently across layers.
+The shallow and middle layers are the most impactful, though jointly tuning all layers produces the strongest results (Sec 3.3).
   3. 3.
-    The impact also varies across attention heads: some heads respond positively to increased attention on this token, while others respond negatively.
-We categorize these as up-effective and down-effective heads, and show that selectively tuning them outperforms uniform tuning across all heads (Sections 2.4 and 2.5).
-These findings motivate ZeroTuning, a simple and effective training-free method that recalibrates the initial token’s attention to improve the LLM’s performance without requiring task-specific token identification (Section  2.6).
-Experiments demonstrate that, despite tuning only one token, ZeroTuning achieves higher average performance on text classification (e.g., SST2, BoolQ), multiple-choice QA (e.g., MMLU, LogiQA), and multi-turn conversation (e.g., MT-Bench) benchmarks across LLMs including Llama-3.1-8B-Instruct, Llama-2-13B-Instruct, Qwen-2-7B, and Deepseek-R1-14B.
-For example, ZeroTuning improves Llama-3.1-8B by 11.71% on classification tasks, 2.64% on QA tasks, and raises its multi-turn score from 7.804 to 7.966.
-The method is also robust under various conditions, including limited resources, few-shot settings, long contexts, quantization, decoding strategies, and prompt variations.
-Our work sheds light on a previously overlooked control point in LLMs, offering new insights into both inference-time tuning and model interpretability.
+    The effect is heterogeneous across attention heads; some respond positively to increased attention (up-effective), while others respond negatively (down-effective).
+We show that selectively targeting the dominant head type outperforms uniform tuning (Sec 3.4,Sec 3.5).
+Building on these findings, we introduce ZeroTuning (see Figure 1), a simple, powerful, and training-free method that recalibrates the initial token’s attention to boost LLM performance without any task-specific identification (Sec 3.6).
+We introduce two variants for attention calibration: a supervised mode by maximizing the accuracy on the labeled validation set, and a novel unsupervised mode by minimizing output entropy.
+Across a suite of 15 benchmarks, ZeroTuning achieves substantial gains on models like Llama-3.1-8B-Instruct, Llama-2-13B-Instruct, Qwen-2-7B, and Deepseek-R1-14B.
+For instance, it boosts Llama-3.1-8B-Instruct performance by a relative 19.9% on classification, 4.5% on question answering, and raises its MT-Bench score from 7.804 to 7.966.
+The method demonstrates remarkable robustness across long contexts, few-shot settings, quantization, and prompt variations.
+Our work not only delivers a practical tool for lightweight model enhancement but also sheds new light on a fundamental control mechanism within LLMs, advancing both inference-time optimization and model interpretability.
+2 Related Work
+Our work is situated at the intersection of two active research areas: inference-time attention tuning and the mechanistic understanding of initial tokens.
+A growing body of work has shown that modifying token-level attention at inference time can enhance the performance of both LLMs and VLMs (Yu et al., 2024; Zhang et al., 2023a; Liu et al., 2024b; Wei & Zhang, 2024).
+However, prevailing methods like PASTA (Zhang et al., 2023a) and Auto-PASTA (Zhang et al., 2024a), which identify and up-weight key tokens, or ACT (Yu et al., 2024), which down-weights non-initial sink tokens, fundamentally rely on heuristics to identify task-specific tokens.
+This reliance limits their universality and introduces potential biases.
+Concurrently, another line of research has focused on explaining why the initial token often becomes an ”attention sink” (Xiao et al., 2023), attributing it to architectural biases and its role as a stabilizing anchor (Barbero et al., 2025; Gu et al., 2024).
+While these studies provide a crucial understanding of what the phenomenon is, the question of how to actively and elegantly harness it for performance gains remains largely unexplored.
+Our work bridges this gap.
+We shift the focus from task-specific token identification to a universal, task-agnostic control point, and move from passive observation of the initial token to a practical tuning framework that leverages its unique properties.
+Detailed related work is provided in Appendix A.
 3 Unveiling the Power of the Initial Token
-In this section, we delve into the impact of tuning the initial token’s attention and introduce our method.
-Unless otherwise specified, all experiments use the Llama-3.1-8B-Instruct model, with dataset details provided in Section 3.1.
+In this section, we first formalize the mechanism of tuning the initial token’s attention, then empirically demonstrate its unique importance and dissect its effects across the model’s layers and heads.
+This systematic analysis culminates in our proposed ZeroTuning methodology.
+Unless otherwise specified, all experiments use the Llama-3.1-8B-Instruct model, with setup details in Section 4.1.
 3.1 Formalizing the Tuning Process
-Figure 1: Effect of the scaling factor $\gamma\in[0,10]$ and initial token’s attention $a_{0}\in\{0,0.2,0.4,0.6,0.8\}$ on attention differences $|a_{i}^{\prime}-a_{j}^{\prime}|$ between non-initial tokens.
-Shaded regions highlight tuning regimes: excessive tuning ($\gamma\in[0,0.3]\cup[6.5,10]$) that may impair LLM performance, moderate sharpening ($\gamma\in[0.3,1]$) that amplifies token-level attention differences, and moderate smoothing ($\gamma\in[1,6.5]$) that reduces them.
-Decreasing $\gamma$ sharpens the distribution by amplifying attention differences, while increasing $\gamma$ smooths it by reducing disparities.
-In decoder-only Transformers, the next token is generated using the final token as the query, attending to all preceding tokens (including itself) as keys.
-Consider a single-layer, single-head attention mechanism with an input sequence:
-$\mathbf{X}=[x_{0},x_{1},\dots,x_{T-1}]\in\mathbb{R}^{d\times T}.$ (1)
-At generation step $T$, the query is $x_{T-1}$, and the attention weights over the keys $x_{0}$ to $x_{T-1}$ form a probability vector:
-$\boldsymbol{a}=[a_{0},a_{1},\dots,a_{T-1}],\quad a_{i}\geq 0,\quad\sum_{i=0}^{T-1}a_{i}=1,$ (2)
-where $a_{0}$ denotes the attention of the initial token, while $a_{1}$ to $a_{T-1}$ correspond to subsequent tokens.
-To control the influence of $x_{0}$, we scale its attention weight by a factor $\gamma>0$ and re-normalize:
+In a decoder-only Transformer, autoregressive generation for a sequence $\mathbf{X}=[x_{0},x_{1},\dots,x_{T-1}]\in\mathbb{R}^{d\times T}$ involves a causal self-attention mechanism.
+At timestep $T$, the query is derived from the final token representation, $x_{T-1}$, which attends to all preceding token representations (including itself) as keys.
+This process yields an attention weight distribution over the input sequence:
+$\bm{a}=[a_{0},a_{1},\dots,a_{T-1}],\quad\text{where}\quad a_{i}\geq 0\quad\text{and}\quad\sum_{i=0}^{T-1}a_{i}=1.$ (1)
+Here, $a_{0}$ is the attention score assigned to the initial token, while $a_{1},\dots,a_{T-1}$ correspond to subsequent tokens.
+To control the influence of $x_{0}$, we introduce a tuning factor $\gamma>0$ to scale its attention and re-normalize:
 $a_{0}^{\prime}=\frac{\gamma a_{0}}{D},\quad a_{i}^{\prime}=\frac{a_{i}}{D}\quad\text{for }i=1,\dots,T-1,$ (2)
-where the normalization constant is
-$D=\gamma a_{0}+\sum_{i=1}^{T-1}a_{i}=\gamma a_{0}+(1-a_{0})=(\gamma-1)a_{0}+1.$ (4)
-This rescaling preserves the relative proportions among non-initial tokens:
+where the normalization constant $D=\gamma a_{0}+\sum_{i=1}^{T-1}a_{i}=(\gamma-1)a_{0}+1$.
+This rescaling preserves the relative proportions among all non-initial tokens:
 $\frac{a_{i}^{\prime}}{\sum_{j=1}^{T-1}a_{j}^{\prime}}=\frac{\frac{a_{i}}{D}}{\sum_{j=1}^{T-1}\frac{a_{j}}{D}}=\frac{a_{i}}{\sum_{j=1}^{T-1}a_{j}},\quad\text{for }i\geq 1,$ (3)
@@
 $a_{i}^{\prime}-a_{j}^{\prime}=\frac{a_{i}-a_{j}}{D}=\frac{a_{i}-a_{j}}{(\gamma-1)a_{0}+1},\quad\text{for }i,j\geq 1.$ (4)
-Intuitively, setting $\gamma>1$ amplifies $a_{0}$ and flattens the attention distribution over the remaining tokens.
-Conversely, $\gamma<1$ suppresses $a_{0}$ and sharpens distinctions among other tokens.
-When $\gamma=1$, the attention remains unchanged.
-As illustrated in Figure 1, the value of $a_{0}$ determines the strength of the tuning effect—larger $a_{0}$ values lead to more pronounced modulation.
-However, extreme values of $\gamma$ can be detrimental: overly small values degrade model performance, a behavior consistent with prior findings on attention sink studies [19], while excessively large values can also hurt performance by suppressing attention to downstream tokens toward near-zero values.
-Then, we analyze how this tuning influences the final representations. Let the value states corresponding to each token be:
-$\mathbf{V}=[\mathbf{v}_{0},\mathbf{v}_{1},\dots,\mathbf{v}_{T-1}]\in\mathbb{R}^{d_{v}\times T},$ (7)
-where $\mathbf{v}_{i}\in\mathbb{R}^{d_{v}}$ is the value vector associated with token $x_{i}$.
-The attention output is a weighted sum of these value vectors.
-Applying the scaled attention weights yields:
-$\mathbf{o}$ $=\sum_{i=0}^{T-1}a_{i}^{\prime}\mathbf{v}_{i}=\frac{\gamma a_{0}}{D}\mathbf{v}_{0}+\sum_{i=1}^{T-1}\frac{a_{i}}{D}\mathbf{v}_{i},$
-$=\frac{1}{D}\left[\gamma a_{0}\mathbf{v}_{0}+\sum_{i=1}^{T-1}a_{i}\mathbf{v}_{i}\right].$ (8)
-Previous studies have observed the “value-state drain” phenomenon, wherein the value state of the initial token tends to be much smaller than those of the other tokens on average [20, 21, 22, 18].
-Therefore, the contribution of the first term $\gamma a_{0}\mathbf{v}_{0}$ to the final output is relatively small, and the value states of the subsequent tokens dominate the final representation.
-Tuning $\gamma$ adjusts the relative contributions of these value states, influencing the final representation.
+Intuitively, $\gamma>1$ amplifies $a_{0}$, flattening the remaining distribution, while $\gamma<1$ suppresses $a_{0}$, sharpening it.
+Theoretically, the magnitude of this effect is governed by the initial token’s own attention weight, $a_{0}$.
+We define this effect, $E_{\text{diff},i,j}$, as the change in attention difference between any two non-initial tokens $i$ and $j$:
+$E_{\text{diff},i,j}=|(a_{i}^{\prime}-a_{j}^{\prime})-(a_{i}-a_{j})|=|a_{i}-a_{j}|\left|\frac{1}{(\gamma-1)a_{0}+1}-1\right|=|a_{i}-a_{j}|\frac{|\gamma-1|a_{0}}{(\gamma-1)a_{0}+1}.$ (5)
+To analyze how $E_{\text{diff},i,j}$ varies with $a_{0}$, we take its partial derivative with respect to $a_{0}$:
+$\frac{\partial E_{\text{diff},i,j}}{\partial a_{0}}=|a_{i}-a_{j}||\gamma-1|\cdot\frac{1}{((\gamma-1)a_{0}+1)^{2}}.$ (6)
+Given that $|a_{i}-a_{j}||\gamma-1|\geq 0$ and the denominator $((\gamma-1)a_{0}+1)^{2}=D^{2}>0$, the derivative is non-negative.
+Thus, $E_{\text{diff},i,j}$ is a monotonically non-decreasing function of $a_{0}$.
+In any non-trivial case ($\gamma\neq 1$ and $a_{i}\neq a_{j}$), the relationship is strictly increasing.
+A detailed proof and visualization are in Appendix C.
+This result provides a crucial insight: the larger the initial token’s attention score ($a_{0}$), the more powerful it becomes as a lever for controlling the entire attention distribution.
+Since prior work has established that initial tokens are natural attention sinks (Barbero et al., 2025), they are inherently potent control points for this tuning process.
 3.2 The Unique Importance of the Initial Token
@@
 Modifying the initial token’s attention consistently yields significant accuracy improvements, often surpassing adjustments to other tokens.
-Given the special role of the initial token in LLMs, we begin by addressing two key questions: (a) Can tuning its attention improve performance across downstream tasks?
+Given the special role of the initial token, we first investigate a key empirical question: (a) Does tuning its attention positively impact performance on downstream tasks?
 (b) Is this position more effective and influential than others?
@@
 Interestingly, performance varies with the direction of tuning: tasks like SST-2 benefit from up-scaling ($\gamma>1$), while BoolQ and LogiQA improve with down-scaling ($\gamma<1$).
-Previous work has identified the initial token as an attention sink that helps prevent over-mixing of information during autoregressive generation [18, 19].
-We extend this understanding by showing that tuning the attention of the initial token can reshape the attention distribution and improve model performance across tasks.
-For example (see Appendix L), in the SST-2 sentiment classification task, input sentences often contain both positive and negative cues.
-We find that LLMs tend to overemphasize isolated negative keywords, leading to biased predictions.
-By increasing attention on the initial token, we flatten the attention distribution, promoting a more balanced integration of information and fostering a more holistic understanding.
-In contrast, in the BoolQ task, where long passages must be inspected for a small number of critical cues, the model may miss relevant tokens due to excessively diffuse attention.
-Reducing initial tokens’ attention amplifies the relative differences between the remaining tokens, sharpening focus on salient input regions and improving response accuracy.
-As discussed in Section 2.1, the higher the attention on the initial token, the greater the impact of such tuning.
-Given that attention sinks are typically concentrated at this position, the initial token is inherently more impactful when modulated, explaining its outsized influence relative to other tokens.
-Besides, unlike other tokens, the initial token is task-agnostic; adjusting its attention does not alter the relative importance of task-relevant tokens, thereby avoiding harmful biases such as inadvertently amplifying attention to irrelevant content.
+Figure 3: The impact of the tuning.
+Previous work has identified the initial token as an attention sink that helps prevent over-mixing of information during autoregressive generation (Gu et al., 2024; Barbero et al., 2025).
+Our empirical results extend this understanding, showing that tuning the initial token’s attention beneficially reshapes the distribution over subsequent tokens and, in turn, improves model performance across tasks.
+This beneficial effect can be understood from two perspectives.
+1.
+Correcting Biases.
+The tuning process counteracts reasoning flaws caused by pretrained attention biases for each dataset, as visualized in Figure 3.
+Our error analysis reveals a clear pattern: (a) Up-scaling ($\gamma>1$) benefits tasks requiring holistic context integration.
+It flattens the attention distribution, preventing the model from over-focusing on misleading local cues.
+For instance, in SST-2 sentiment analysis (see Appendix F), LLMs often fixate on isolated negative keywords while ignoring the surrounding positive context.
+Increasing the initial token’s attention promotes a more balanced, global understanding, thereby correcting such biased predictions.
+(b) Down-scaling ($\gamma<1$) benefits tasks demanding sharp focus on critical evidence.
+It amplifies the relative weights of non-initial tokens, helping the model pinpoint key details in a diffuse context.
+In long-context tasks like BoolQ, a model’s focus can become diluted, failing to locate the precise text segment containing the answer.
+Reducing the initial token’s attention sharpens the model’s focus on this salient information.
+Figure 4: A strong inverse correlation between accuracy (orange) and next-token prediction entropy (blue).
+2.
+Reducing Predictive Uncertainty.
+The tuning process can be viewed through the lens of output entropy, a proxy for model uncertainty.
+As illustrated in Figure 4, a clear inverse correlation emerges: the scaling factor that minimizes entropy consistently aligns with the factor that maximizes accuracy.
+This suggests our method better unlocks the model’s pretrained knowledge, leading to more confident and correct predictions.
 3.3 Layer-wise Analysis of Initial Token Scaling
-To better understand the impact of tuning the initial token’s attention, we examine how its effect varies when applied selectively across different layers.
-Following prior work on layer functionality in transformer-based models [23, 24], we divide the 32 layers of Llama-3.1-8B-Instruct into three groups: shallow layers (Layers 1–10), middle layers (Layers 11–21), and deep layers (Layers 22–31).
+To understand the propagation of this effect, we examine how its effect varies when applied selectively across different layers.
+Following prior work on layer functionality in transformer-based models (Jin et al., 2024; Zhang et al., 2024b), we divide the 32 layers of Llama-3.1-8B-Instruct into three groups: shallow (Layers 1–10), middle (Layers 11–21), and deep (Layers 22–31).
 We then perform independent attention scaling experiments for each group on six tasks: BoolQ, SST-2, SST-5, MR, LogiQA, and MathQA.
@@
 Different depths exhibit a consistent accuracy trend with varying magnitudes.
-As shown in Figure 3, tuning the initial token’s attention yields consistent trends across all three layer groups (i.e., accuracy changes similarly with scaling regardless of depth), and jointly tuning all layers amplifies these benefits, often resulting in the highest accuracy.
+As shown in Figure 5, tuning yields consistent trends at any depth (i.e., accuracy changes similarly with scaling regardless of depth), and jointly tuning all layers amplifies these benefits, often resulting in the highest accuracy.
 However, the magnitude of improvement varies.
 In most cases, tuning the shallow and middle layers leads to greater accuracy than tuning the deep layers.
-This aligns with prior findings that early and middle layers primarily contribute to representation learning and knowledge integration, whereas deep layers focus on task-specific reasoning over aggregated features [25, 26].
+Prior studies have found that early and middle layers mainly support representation learning and knowledge integration, while deep layers focus on task-specific reasoning over aggregated features (Chen et al., 2024; Jin et al., 2025).
+Therefore, we argue that the tuning process more effectively reshapes the representational space in shallow and middle layers, promoting better downstream performance and reducing uncertainty.
 3.4 Analyzing the Role of the Initial Token Across Attention Heads
@@
 Results reveal heterogeneous behavior among heads, motivating head-specific tuning strategies.
-Unlike layers, which pass information sequentially, attention heads operate in parallel and contribute independently via concatenation.
+Unlike layers passing information sequentially, attention heads operate in parallel and contribute independently via concatenation.
 It remains unclear how they differ in response to the initial token.
@@
 In contrast, MMLU has a higher proportion of down-effective heads, making down-scaling more effective than up-scaling.
-These findings align with prior work suggesting that attention heads specialize into distinct functional roles during pretraining [27].
-Such roles include global retrieval, structural parsing, option discrimination, and sensitivity to negation.
-We hypothesize that these functional differences drive the varying impact of initial token attention scaling.
-While some heads rely on broad attention to support global reasoning, others benefit from focused attention on salient tokens.
-Together, these specialized mechanisms enable language models to better understand and process diverse input patterns.
+These results align with prior studies showing that attention heads specialize into distinct functional roles during pretraining (Zheng et al., 2024b; Guo et al., 2024), such as global retrieval, structural parsing, option discrimination, and negation sensitivity.
+We propose that these functional differences may explain the variable impact of initial token attention scaling, with some heads supporting broad global reasoning and others focusing on salient tokens.
+This interpretation requires further exploration in future work.
 3.5 Evaluating Head-Specific Tuning Strategies
-Figure 5: Accuracy comparison of different tuning strategies on (a) SST-2 and (b) MMLU.
-Head-specific tuning (Up, Down) consistently outperforms uniform scaling, validating the importance of accounting for head-level behavior.
 Given the diversity in head responses, we investigate whether head-specific tuning offers greater effectiveness than uniform tuning.
@@
 Interestingly, the Up+Down strategy does not outperform Up or Down individually, possibly due to the concatenative nature of attention heads and suboptimal joint scaling.
+Figure 7: Accuracy comparison of different tuning strategies on (a) SST-2 and (b) MMLU.
+Head-specific tuning (Up, Down) consistently outperforms uniform scaling, validating the importance of accounting for head-level behavior.
 3.6 ZeroTuning Methodology
-Building on our empirical findings, we propose ZeroTuning, a simple yet effective method that adjusts the initial token’s attention in a head-specific manner to enhance LLM performance, without requiring task-specific token identification.
-ZeroTuning comprises three steps:
+Building on our empirical findings, we propose ZeroTuning, a method that enhances LLM performance via head-specific attention adjustments to the initial token, without requiring task-specific token identification.
+The core methodology involves three steps: head behavior profiling, selective rescaling, and renormalization.
+To accomplish the first two steps, hyperparameter calibration, we introduce both supervised and unsupervised variants to cater to different application scenarios.
+Supervised Calibration
+Consistent with standard practices in inference-time adaptation (Yu et al., 2024; Zhang et al., 2023a; 2024a), our primary approach utilizes a labeled calibration set (e.g., validation examples) to perform:
   1. 1.
-    Head Behavior Profiling: Using a small calibration set (e.g., a few validation examples), we assess each attention head’s sensitivity to scaling the initial token’s attention.
-A head is classified as up-effective if increased attention improves performance, and down-effective otherwise.
+    Head Behavior Profiling: Assessing each attention head’s sensitivity to the initial token’s attention scaling.
+A head is classified as up-effective if increased attention improves accuracy, and down-effective otherwise.
   2. 2.
-    Selective Rescaling: Based on the dominant head type (i.e., the group with the highest proportion), we apply a scaling factor $\gamma$ exclusively to that group to adjust the initial token’s attention.
-[^1] We explore the effect of scaling different subsets of heads; results (Figure 8, Appendix F) show that tuning the top 20% to 70% yields better performance.
-  3. 3.
-    Renormalization: The scaled attention scores are re-normalized using the softmax function to maintain a valid distribution.
-For optimized attention implementations (e.g., SDPA, Flash Attention) where direct modification of attention scores is infeasible, ZeroTuning applies scaling to the query or key states before attention computation.
-We demonstrate that tuning key states yields similar effects to attention-based tuning, but induces steeper changes in the attention distribution of subsequent tokens (see Appendix A).
-The following sections evaluate the performance and robustness of ZeroTuning across different tasks and model settings.
+    Selective Rescaling: Applying a scaling factor $\gamma$, identified by searching for the value that maximizes accuracy on the calibration set, exclusively to the dominant head type (i.e., the most numerous group).
+Unsupervised Calibration via Entropy Minimization
+To mitigate the reliance on labeled data, we propose a novel unsupervised calibration strategy.
+Based on our finding that a model’s output entropy strongly correlates with its accuracy (as detailed in Section 3.2), this variant bypasses the calibration set entirely.
+Instead, it identifies the optimal heads and scaling factor $\gamma$ by minimizing the average next-token prediction entropy directly on the unlabeled test inputs.
+Finally, the third step, Renormalization, is applied in both variants by re-normalizing the scaled attention scores via the softmax function to maintain a valid probability distribution.
+For optimized attention implementations (e.g., SDPA, Flash Attention) where direct score modification is infeasible, ZeroTuning applies scaling to the query or key states.
+We demonstrate and show that this yields similar effects in Appendix G.
 4 Experimental Results
 4.1 Experimental Setup
-Models, Tasks, and Datasets.
-Models: We evaluate ZeroTuning on four LLMs with distinct attention implementations: Llama-3.1-8B-Instruct [28] and Llama-2-13B-Chat [29] with eager attention, Qwen-2-7B [30] with SDPA attention, and DeepSeek-R1-14B [31] with Flash attention.
-[^2] Eager, SDPA, and Flash are official attention implementations in modern Transformer libraries.
-Eager computes the full attention map; SDPA uses PyTorch’s efficient API to select the optimal implementation; Flash relies on fused CUDA kernels from the FlashAttention library.
-Tasks and Datasets: Our experiments encompass three task types across 15 datasets: (1) Text Classification and Reasoning, including SST-2 (binary sentiment classification) [32], SST-5 (fine-grained sentiment analysis) [32], MR (movie review polarity detection) [33], SUBJ (subjectivity classification) [34], TREC (question type classification) [35], CB (commitment detection) [36], and BoolQ (boolean question answering) [37]; (2) Domain-Specific Multiple-Choice, including MMLU (cross-domain knowledge testing) [38], AQUA (math word problems) [39], MathQA (algebraic reasoning) [40], LogiQA (logical reasoning) [41], CQA (commonsense reasoning) [42], PIQA (physical commonsense QA) [43], and ARCC (scientific reasoning) [44]; and (3) Multi-Round Conversation, using MT-Bench [39].
-Baselines and Evaluation Metrics.
-Baselines: We benchmark ZeroTuning against three baselines: (1) vanilla inference, which performs standard inference without any modifications; (2) ACT [12], which identifies sink tokens using an attention score threshold and reduces their attention weights; and (3) Auto-PASTA [10], which leverages an LLM to locate important tokens and enhance their attention weights.
-For ACT, we use the original hyperparameters ($\beta=0.4$) and filter attention heads on a 10% validation set.
-For AutoPASTA, the evaluated LLM identifies key input regions based on prompts provided in Appendix J.
-Since ACT relies on explicit attention maps, it is only evaluated on LLaMA-3.1-8B-Instruct with eager attention implementation.
-Evaluation Metrics: We assess performance using accuracy for text classification and multiple-choice tasks.
-For the multi-round conversation task, we report average quality scores as evaluated by GPT-4, following the methodology outlined in [39].
-Implementation Details.
-ZeroTuning is implemented using PyTorch and the Hugging Face Transformers library.
-We select zero-shot setting with standard greedy decoding for baseline consistency, and use prompt templates from [45, 46, 47] across datasets (see Appendix J for details).
-Unless otherwise specified, we tune the top 40% of the heads.
-Each experiment uses a randomly selected validation set of 500 samples with the random seed fixed to 42.
-Table 1: Performance Comparison of Classification Tasks Across Models. The best performance in each dataset is bolded, the second best is underlined, and the ZeroTuning method is highlighted in gray.
-Model Method Datasets SST-2 SST-2 MR BoolQ CB TREC SUBJ Avg.
-Llama-3.1-8B-Instruct Vanilla 73.20 45.40 89.20 69.60 82.14 14.00 44.60 59.59 ACT 85.00 43.80 90.80 58.60 82.14 15.80 44.60 60.11 Auto-PASTA 89.60 47.20 91.40 72.60 83.93 16.00 45.40 63.73 ZeroTuning 91.60 52.00 92.00 82.40 89.29 26.20 66.60 71.44 Qwen-2-7B (SDPA) Vanilla 78.80 45.40 72.40 85.00 78.50 12.60 13.00 55.10 ACT / / / / / / / / Auto-PASTA 89.00 47.00 77.70 85.00 89.29 14.00 57.00 65.57 ZeroTuning 89.60 47.20 87.40 86.40 85.71 26.60 54.40 68.19 Deepseek-R1-14B (Flash) Vanilla 91.20 49.40 89.20 83.40 89.29 20.80 50.40 67.67 ACT / / / / / / / / Auto-PASTA 92.00 52.20 89.80 83.40 92.86 22.60 50.40 69.04 ZeroTuning 93.00 51.20 90.20 88.00 92.86 32.00 55.80 71.87
-Table 2: Performance Comparison of Multiple-Choice Tasks Across Models
+Our evaluation includes four recent LLMs with distinct attention implementations (Llama-3.1-8B, Llama-2-13B, Qwen-2-7B, and DeepSeek-R1-14B).
+We test performance across 15 datasets spanning three categories: Text Classification, Multiple-Choice QA, and Multi-Round Conversation.
+We benchmark ZeroTuning against three methods: (1) Vanilla inference; (2) ACT (Yu et al., 2024), a sink-token down-scaling method  [^1] Since ACT explicitly manipulates attention maps, we only evaluate it on Llama-3.1-8B-Instruct; and (3) Auto-PASTA (Zhang et al., 2024a), an LLM-guided key-token up-scaling method.
+All experiments are conducted in a zero-shot setting with greedy decoding for fair comparison.
+For our supervised variant and the baselines, hyperparameters are calibrated on a fixed validation set.
+A detailed description of all models, datasets, baselines, and implementation specifics is provided in Appendix B.
+Table 1: Performance Comparison of Classification Tasks Across Models. The best performance in each dataset is bolded and the ZeroTuning method is highlighted in gray.
+Model Method Datasets
+SST2 SST5 MR BoolQ CB TREC SUBJ Avg.
+Llama-3.1-8B-Instruct Vanilla 73.20 45.40 89.20 69.60 82.14 14.00 44.60 59.59
+ACT 85.00 43.80 90.80 58.60 82.14 15.80 44.60 60.11
+Auto-PASTA 89.60 47.20 91.40 72.60 83.93 16.00 45.40 63.73
+[…]
+Qwen-2-7B (SDPA) Vanilla 78.80 45.40 72.40 85.00 78.50 12.60 13.00 55.10
+ACT
+Auto-PASTA 89.00 47.00 77.70 85.00 89.29 14.00 57.00 65.57
+[…]
+Deepseek-R1-14B (Flash) Vanilla 91.20 49.40 89.20 83.40 89.29 20.80 50.40 67.67
+ACT
+Auto-PASTA 92.00 52.20 89.80 83.40 92.86 22.60 50.40 69.04
+[…]
+Table 2: Performance Comparison of Multiple-Choice Tasks Across Models.
 Model Method Datasets
@@
 Auto-PASTA 67.00 31.23 35.20 40.40 78.20 84.60 84.62 60.18
-ZeroTuning 68.80 30.43 36.60 42.80 80.40 85.40 85.95 61.48
+[…]
 Qwen-2-7B (SDPA) Vanilla 69.80 36.76 39.20 45.00 78.80 85.20 86.96 63.10
-ACT / / / / / / / /
+ACT
 Auto-PASTA 69.80 39.13 39.20 45.00 82.60 85.40 86.96 64.01
-ZeroTuning 70.40 39.92 40.20 47.40 81.80 86.20 87.96 64.84
+[…]
 Deepseek-R1-14B (Flash) Vanilla 66.60 38.74 38.20 27.80 78.20 84.20 86.62 60.05
-ACT / / / / / / / /
+ACT
 Auto-PASTA 66.60 38.74 39.40 28.20 78.20 84.40 86.62 60.31
-ZeroTuning 70.00 39.13 39.80 35.60 78.60 85.00 87.29 62.20
+[…]
 4.2 Overall performance of ZeroTuning
+For a fair and direct comparison with existing supervised baselines, our main experiments focus on the supervised ZeroTuning approach.
 Text Classification
@@
 Despite tuning only a single token, ZeroTuning consistently outperforms baselines and methods that require tuning more tokens.
-With Llama-3.1-8B-Instruct, it achieves an average improvement of 11.71% over vanilla, with peaks of 22.00% on SUBJ and 18.40% on SST-2.
+With Llama-3.1-8B-Instruct, it achieves an average improvement of +11.71% over vanilla, with peaks of +22.00% on SUBJ and +18.40% on SST-2.
 It outperforms AutoPASTA by an average of 7.71%.
-On Qwen-2-7B, ZeroTuning gains 13.09%, and on Deepseek-R1-14B, it improves by 4.20%, with a notable increase of 11.20% on TREC.
+On Qwen-2-7B, ZeroTuning gains +13.09%, and on Deepseek-R1-14B, it improves by +4.20%, with a notable increase of +11.20% on TREC.
 Domain-Specific Multiple Choice
 Next, we evaluate ZeroTuning on common domain-specific multiple-choice datasets under various settings, as shown in Table 2.
-For Llama-3.1-8B-Instruct, it increases the average accuracy by 2.64%, with gains of 3.40% on LogiQA and 1.40% on MMLU.
-Qwen-2-7B gains 1.74%, and Deepseek-R1-14B gains 2.15%, with an outstanding 7.80% on LogiQA.
+For Llama-3.1-8B-Instruct, it increases the average accuracy by +2.64%, with gains of +3.40% on LogiQA and +1.40% on MMLU.
+Qwen-2-7B gains +1.74%, and Deepseek-R1-14B gains +2.15%, with an outstanding +7.80% on LogiQA.
 Multi-Round Conversation
-We further demonstrate ZeroTuning’s effectiveness in multi-round conversations using MT-Bench [48], with results in Table 3.
+We further demonstrate ZeroTuning’s effectiveness in multi-round conversations using MT-Bench (Zheng et al., 2024a), with results in Table 3.
 For Llama-3.1-8B-Instruct, ZeroTuning improves the average score by 0.162 points (7.966 vs.
@@
 Llama-2-13B-Chat-vanilla 7.063 6.238 6.650
-4 Further Analysis
-We provide additional analyses to deepen the understanding of ZeroTuning’s behavior and practical utility under varying conditions.
-(1) Tuning Different Matrices: Appendix A begins with a comparison between tuning attention scores and key states.
-Tuning attention scores yields stable improvements due to its linear influence, whereas tuning key states is sensitive and can cause sharp performance drops because of exponential effects.
-(2) Resource Constraints: Appendix B evaluates ZeroTuning across three simulated computational budgets, revealing consistent gains from minimal scaling (Level 0) to full head classification and parameter search (Level 2).
-(3) Context Length Sensitivity: Appendix C tests robustness to long contexts with irrelevant tokens, showing ZeroTuning maintains accuracy where vanilla models degrade.
-(4) Few-Shot Learning Robustness: Appendix D demonstrates stable few-shot gains and fewer invalid outputs, indicating ZeroTuning complements in-context learning.
-(5) Prompt Variation Sensitivity: Appendix G assesses prompt modifications such as dropped or altered instructions; ZeroTuning sustains strong performance.
-(6) Effect of Attention Head Subset Selection: Appendix F shows that tuning a moderate proportion of attention heads (20%–70%) outperforms both full and minimal tuning.
-(7) Quantization Robustness: Appendix H shows a consistent trend across 16-, 8-, and 4-bit quantization, with ZeroTuning partially mitigating degradation from low-precision quantization.
-(8) Decoding Strategy Impact: Appendix E evaluates various decoding methods (Top-$k$, Top-$p$, Beam Search), confirming ZeroTuning’s effectiveness.
-5 Related Work
-5.1 Token-Level Attention Tuning
+4.3 Unsupervised ZeroTuning
+We validate the fully unsupervised variant of ZeroTuning, which eliminates the need for any labeled calibration data by minimizing the model’s average output entropy.
+Appendix E provides a detailed empirical validation for this approach, including visual comparisons, empirical analysis, and a breakdown of error patterns.
+As shown in Figure 8, this entropy-guided method achieves performance highly competitive with its supervised counterpart, extending its applicability to label-scarce scenarios.
+Figure 8: Performance comparison of Vanilla, Supervised, and Unsupervised ZeroTuning on Llama-3.1-8B-Instruct.
+5 Further Analysis and Key Findings
+Robustness Across Diverse Conditions.
+ZeroTuning demonstrates remarkable stability.
+It maintains strong performance gains even when faced with: (1) long contexts padded with irrelevant distractors, where it stabilizes information flow better than the vanilla model (Appendix I); (2) few-shot scenarios, where it consistently improves instruction-following and reduces invalid outputs (Appendix J); (3) significant prompt variations, including missing or altered instructions (Appendix M); and (4) low-precision 4-bit and 8-bit quantization, where it partially mitigates the associated accuracy degradation (Appendix N).
+Practicality and Methodological Choices.
+Our method is not only robust but also highly practical.
+We theoretically and empirically confirm that tuning key states provides a viable, kernel-agnostic alternative to direct attention score manipulation, proving effective in optimized environments like FlashAttention (Appendix G).
+Furthermore, ZeroTuning is adaptable to resource-constrained settings, delivering gains even with a minimal, search-free scaling approach (Appendix H).
+We also analyze key methodological choices, showing that tuning a moderate subset of heads (40%–70%) is optimal, providing a clear and efficient configuration (Appendix L).
+Boundaries of Efficacy.
+Finally, we analyze the method’s boundaries and potential side effects.
+We quantitatively demonstrate that ZeroTuning excels at correcting a model’s uncertain errors but cannot override high-confidence mistakes rooted in flawed pretrained knowledge.
+We also characterize the negative effects of extreme tuning, which provides a clearer picture of the method’s operational limits (Appendix D).
+This positions our method as a powerful tool for unlocking a model’s latent knowledge, rather than a substitute for fine-tuning.
+Intriguingly, we also find that within a safe operational range, the scaling factor can modulate output diversity in a manner analogous to temperature, but with the unique ability to alter the rank-ordering of logits and thereby correct errors that temperature scaling cannot fix.
+6 Conclusion
+In this work, we present a systematic analysis of tuning the initial token’s attention and propose ZeroTuning, a novel, training-free method to enhance LLMs.
+By recalibrating this single, task-agnostic token, ZeroTuning outperforms previous methods that require task-specific tuning.
+It operates effectively in both supervised and unsupervised modes and demonstrates broad compatibility across various implementations.
+This work advances inference-time tuning and contributes to the interpretability of LLMs, opening new avenues for lightweight model optimization.
+Acknowledgments
+We extend special thanks to Prof. Surbhi Goel from the University of Pennsylvania for her valuable advice on the theoretical aspects of this research.
+Appendix A Related Work
+A.1 Token-Level Attention Tuning
 Token-level attention tuning typically aims to increase attention to critical input tokens or decrease attention to less informative tokens.
-[49] proposes a mask perturbation method to adjust attention weights for key tokens, thereby improving translation quality.
-[11] introduce PASTA, which allows manual designation of important tokens during inference.
-This is extended by AutoPASTA [10], which uses LLMs to autonomously identify salient tokens and increase attention to them.
-In contrast, ACT [12] reduces attention to semantically trivial sink tokens and redirects it to meaningful content.
+Lu et al.
+(2021) proposes a mask perturbation method to adjust attention weights for key tokens, thereby improving translation quality.
+Zhang et al.
+(2023a) introduces PASTA, which allows manual designation of important tokens during inference.
+This is extended by AutoPASTA (Zhang et al., 2024a), which uses LLMs to autonomously identify salient tokens and increase attention to them.
+In contrast, ACT (Yu et al., 2024) reduces attention to semantically trivial sink tokens and redirects it to meaningful content.
 Similar strategies have been applied to VLMs to mitigate hallucinations.
-PAI [13] enhances attention to image tokens at inference time to counteract text-dominant bias.
-IBD [14] and OPERA [15] further refine this idea by prioritizing visual information or penalizing overconfident summary tokens.
+PAI (Liu et al., 2024b) enhances attention to image tokens at inference time to counteract text-dominant bias.
+IBD (Zhu et al., 2024) and OPERA (Wei & Zhang, 2024) further refine this idea by prioritizing visual information or penalizing overconfident summary tokens.
 While effective, these methods depend on identifying task-specific tokens, which may introduce bias (e.g., overemphasizing misleading tokens) and limit applicability when token importance is unclear or attention maps are unavailable.
 In contrast, our method focuses on a task-invariant initial token, removing the need for costly token identification, and can be easily applied by tuning key states.
-5.2 The Magic of the Initial Token
+A.2 The Magic of the Initial Token
 Recent studies highlight the significance of the initial token, especially through the lens of the attention sink phenomenon, where it draws substantial attention despite low semantic content.
-[16] show that preserving such tokens is critical for maintaining performance in sliding window attention.
-[17] attribute this effect to softmax normalization and causal masking, while [18] and [19] identify architectural biases that amplify attention to the initial token, including key-query alignment and LayerNorm effects.
-Functionally, the initial token is hypothesized to serve as a stabilizing “no-op” anchor, enhancing robustness to prompt variations [19].
-It has been leveraged in applications such as long-context modeling [50, 16], but also poses challenges for quantization due to its high attention weight [2, 51].
+Xiao et al.
+(2023) shows that preserving such tokens is critical for maintaining performance in sliding window attention.
+Kaul et al.
+(2024) attributes this effect to softmax normalization and causal masking, while Gu et al.
+(2024) and Barbero et al.
+(2025) identify architectural biases that amplify attention to the initial token, including key-query alignment and LayerNorm effects.
+Functionally, the initial token is hypothesized to serve as a stabilizing “no-op” anchor, enhancing robustness to prompt variations (Barbero et al., 2025).
+It has been leveraged in applications such as long-context modeling (Zhang et al., 2023b; Xiao et al., 2023), but also poses challenges for quantization due to its high attention weight (Dettmers et al., 2023; Liu et al., 2024a).
 While previous work has identified the structural and functional importance of the initial token, its potential as a target for attention tuning remains underexplored.
 In this work, we provide a detailed analysis of attention tuning of the initial token across layers and heads, demonstrating its consistent influence across different tasks.
-Our approach bridges the gap between the these lines of research by proposing a novel method that advances interpretable attention tuning.
-6 Conclusion
-In this work, we propose ZeroTuning, a novel training-free method to enhance large language models (LLMs) by adjusting the attention of the initial token.
-Our findings reveal that tuning the initial token’s attention can lead to substantial improvements in LLM performance across a variety of tasks, including text classification, question answering, and multi-round conversations.
-Notably, ZeroTuning outperforms previous task-specific tuning methods, achieving robust and transferable gains without requiring task-specific token identification.
-The approach offers a fresh perspective on optimizing LLMs by leveraging the often-overlooked power of the initial token, making it a simple yet effective solution for enhancing model performance.
-This work advances inference-time tuning and contributes to the interpretability of LLMs.
-It opens up new avenues for further exploration and optimization in the field.
-Appendix A The Effect of Tuning Different Matrices
+Our approach bridges the gap between these lines of research by proposing a novel method that advances interpretable attention tuning.
+Appendix B Detailed Experimental Setup
+Models, Tasks, and Datasets.
+Models: We evaluate ZeroTuning on four LLMs with distinct attention implementations: Llama-3.1-8B-Instruct (Grattafiori et al., 2024) and Llama-2-13B-Chat (Touvron et al., 2023) with eager attention, Qwen-2-7B (Yang et al., 2024) with SDPA attention, and DeepSeek-R1-14B (DeepSeek-AI et al., 2025) with Flash attention.
+[^2] Eager, SDPA, and Flash are official attention implementations in modern Transformer libraries.
+Eager computes the full attention map; SDPA uses PyTorch’s efficient API to select the optimal implementation; Flash relies on fused CUDA kernels from the FlashAttention library.
+Tasks and Datasets: Our experiments encompass three task types across 15 datasets: (1) Text Classification and Reasoning, including SST-2 (binary sentiment classification) (Socher et al., 2013), SST-5 (fine-grained sentiment analysis) (Socher et al., 2013), MR (movie review polarity detection) (Pang & Lee, 2005), SUBJ (subjectivity classification) (Pang & Lee, 2004), TREC (question type classification) (Li & Roth, 2002), CB (commitment detection) (De Marneffe et al., 2019), and BoolQ (boolean question answering) (Clark et al., 2019); (2) Domain-Specific Multiple-Choice, including MMLU (cross-domain knowledge testing) (Hendrycks et al., 2020), AQUA (math word problems) (Zheng et al., 2024a), MathQA (algebraic reasoning) (Amini et al., 2019), LogiQA (logical reasoning) (Liu et al., 2023), CQA (commonsense reasoning) (Talmor et al., 2018), PIQA (physical commonsense QA) (Bisk et al., 2020), and ARCC (scientific reasoning) (Clark et al., 2018); and (3) Multi-Round Conversation, using MT-Bench (Zheng et al., 2024a).
+Baselines and Evaluation Metrics.
+Baselines: We benchmark ZeroTuning against three baselines: (1) vanilla inference, which performs standard inference without any modifications; (2) ACT (Yu et al., 2024), which identifies none-initial sink tokens using an attention score threshold and reduces their attention weights; and (3) Auto-PASTA (Zhang et al., 2024a), which leverages an LLM to locate important tokens and enhance their attention weights.
+Evaluation Metrics: We assess performance using accuracy for text classification and multiple-choice tasks.
+For the multi-round conversation task, we report average quality scores as evaluated by GPT-4, following the methodology outlined in Zheng et al.
+(2024a).
+Implementation Details.
+All experiments are implemented in PyTorch using the Hugging Face Transformers library.
+We use a zero-shot setting with greedy decoding for consistency across all methods.
+For our supervised variant and the baselines, we use a fixed validation set of 500 randomly selected samples (seed 42) for calibration.
+For ZeroTuning, we tune the top 40% of identified heads unless otherwise specified.
+For ACT, we use the official hyperparameter ($\beta=0.4$), and since it requires explicit attention maps, we only evaluate it on Llama-3.1-8B-Instruct.
+Prompts for all tasks and baselines are detailed in Appendix O.
+Appendix C Theoretical Analysis of Tuning Efficacy via the Initial Token
+This appendix provides a formal proof for the claim made in Section 3.1: that the tuning effect’s magnitude is governed by the initial token’s attention weight $a_{0}$.
+Proposition 1.
+For any given scaling factor $\gamma\neq 1$ and any two non-initial tokens $i,j\geq 1$ with unequal initial attention weights ($a_{i}\neq a_{j}$), the magnitude of the tuning effect on their attention difference is a strictly monotonically increasing function of the initial token’s attention weight, $a_{0}$.
+We aim to show that the tuning effect, $E_{\text{diff},i,j}$, is a monotonically increasing function of $a_{0}$ by proving its partial derivative with respect to $a_{0}$ is positive.
+Recall the definition of the effect magnitude from Eq.
+equation 5:
+$E_{\text{diff},i,j}(a_{0})=|a_{i}-a_{j}|\frac{|\gamma-1|a_{0}}{(\gamma-1)a_{0}+1}.$ (7)
+Taking the partial derivative of $E_{\text{diff}}$ with respect to $a_{0}$, we treat the term $|a_{i}-a_{j}||\gamma-1|$ as a constant factor:
+$\frac{\partial E_{\text{diff},i,j}}{\partial a_{0}}$ $=|a_{i}-a_{j}||\gamma-1|\cdot\frac{\partial}{\partial a_{0}}\left(\frac{a_{0}}{(\gamma-1)a_{0}+1}\right)$ (8)
+$=|a_{i}-a_{j}||\gamma-1|\cdot\frac{1\cdot((\gamma-1)a_{0}+1)-a_{0}\cdot(\gamma-1)}{((\gamma-1)a_{0}+1)^{2}}$ (9)
+$=|a_{i}-a_{j}||\gamma-1|\cdot\frac{1}{((\gamma-1)a_{0}+1)^{2}}.$ (10)
+The term $|a_{i}-a_{j}||\gamma-1|$ is non-negative.
+The denominator, $((\gamma-1)a_{0}+1)^{2}=D^{2}$, is the square of the normalization constant and is strictly positive for any valid probability distribution.
+Therefore, the derivative $\frac{\partial E_{\text{diff},i,j}}{\partial a_{0}}\geq 0$.
+Furthermore, for any non-trivial case where the tuning factor is active ($\gamma\neq 1$) and the attention weights are not uniform ($a_{i}\neq a_{j}$ for some $i,j$), the derivative is strictly positive.
+This proves that $E_{\text{diff}}$ is a strictly monotonically increasing function of $a_{0}$.
+Consequently, a larger initial attention weight provides a more powerful lever for modulating the attention distribution.
+This theoretical result is visually corroborated by Figure 9.
+Figure 9: Visualization of the tuning effect on attention differences $|a_{i}^{\prime}-a_{j}^{\prime}|$ as a function of the scaling factor $\gamma$ for various initial attention weights $a_{0}\in\{0,0.2,0.4,0.6,0.8\}$.
+The plot demonstrates that a higher initial attention weight $a_{0}$ (e.g., the red curve) leads to a significantly stronger response to changes in $\gamma$.
+We identify three primary regimes: sharpening ($\gamma<1$), where attention differences are amplified; smoothing ($\gamma>1$), where differences are diminished; and regions of excessive tuning (e.g., $\gamma\to 0$ or $\gamma\gg 1$), which may degrade performance.
+Appendix D Deeper Analysis of Potential Failures and Negative Effects
+To provide a comprehensive understanding of ZeroTuning, we analyze its operational boundaries and potential negative side effects when pushed to its limits.
+Boundaries of Efficacy: Unlocking Latent Knowledge vs. Correcting Factual Errors.
+Our analysis reveals a key insight into ZeroTuning’s mechanism: it primarily unlocks and disambiguates a model’s latent knowledge, rather than correcting deeply ingrained factual errors.
+To test this, we quantitatively analyzed ZeroTuning’s corrective power as a function of the vanilla model’s initial prediction confidence.
+We partitioned the set of SST2 incorrect predictions into ”uncertain errors” (vanilla softmax confidence << threshold) and ”certain errors” (confidence $\geq$ threshold) and evaluated our method’s performance on each group.
+Figure 10: Corrected accuracy on initially incorrect samples as a function of the vanilla model’s confidence (certainty threshold).
+The results, shown in Figure 10, provide strong empirical validation.
+ZeroTuning is highly effective on the uncertain error population (blue line), correcting over 80% of mistakes where the vanilla model’s confidence was below 0.5.
+Conversely, its ability to fix certain errors (red line) decays sharply as the base model’s confidence increases, dropping to near-zero on predictions where the model was already confidently wrong.
+This confirms a clear operational boundary: ZeroTuning excels at resolving low-confidence mistakes by refining the model’s focus, but it is not designed to overwrite high-confidence knowledge learned during pretraining.
+This finding positions ZeroTuning not as a replacement for fine-tuning, but as a powerful, complementary inference-time technique.
+It assists a model in better leveraging its existing, albeit sometimes uncertain, knowledge.
+The potential for synergistic interaction between ZeroTuning and parameter-efficient fine-tuning methods like LoRA remains a promising avenue for future research.
+Negative Effects of Extreme Tuning.
+We also investigated the effects of applying extreme scaling factors ($\gamma$), far outside the optimal range.
+These experiments reveal predictable failure modes that further illuminate the role of the initial token:
+  • Overly-suppressed attention ($\gamma\to 0$): When the initial token’s attention is excessively reduced, the model’s output often becomes degenerative.
+We observe a tendency for the model to enter repetitive loops, outputting a single answer (e.g., ”True”) without any of the semantic elaboration or reasoning present in the vanilla output.
+This suggests that a minimal level of attention to the ”sink” token is necessary to maintain generative stability.
+  • Overly-amplified attention ($\gamma\gg 1$): Conversely, when the initial token’s attention is excessively high, it can disrupt the model’s ability to follow complex instructions.
+By absorbing too much of the attention budget, the initial token appears to prevent other, more task-relevant tokens from receiving the focus they need, leading to incomplete or non-compliant answers.
+Interestingly, within a reasonable range, moderate tuning of the initial token’s attention can produce effects analogous to adjusting the temperature parameter in decoding.
+It can modulate the diversity of the output, encouraging the model to explore different perspectives or generate more varied responses, without the repetitive downsides of extreme scaling.
+However, our method is fundamentally more powerful for error correction.
+Temperature scaling acts on the final logits $z$ just before the softmax, calculating the probability of the $i$-th token as $p_{i}=\text{softmax}(z_{i}/T)$.
+Since dividing by a positive temperature $T$ does not change the relative order of the logits (i.e., $\arg\max_{i}(z_{i})=\arg\max_{i}(z_{i}/T)$), temperature scaling cannot alter the outcome of greedy decoding.
+In contrast, ZeroTuning operates at the attention level, optimizing the model’s internal representations.
+This process produces an entirely new set of output logits, $z^{\prime}$, which can have a different rank ordering.
+It is therefore possible for the originally predicted token $\arg\max_{i}(z_{i})$ to be incorrect, while the new prediction $\arg\max_{i}(z^{\prime}_{i})$ becomes correct, enabling error correction.
+Appendix E Unsupervised ZeroTuning: Analysis and Results
+Figure 11: Visualizing the relationship between the scaling factor $\gamma$ and three key metrics: average entropy (left), log-entropy (middle), and task accuracy (right).
+Across diverse datasets, the entropy minimum consistently aligns with a region of high accuracy, validating entropy as a strong signal for unsupervised tuning.
+This section details the unsupervised variant of ZeroTuning, which eliminates the need for a labeled calibration set by leveraging the model’s output entropy as a proxy for performance.
+We begin with a visual analysis to establish the core principle behind this approach.
+Figure 11 plots three key metrics against the attention scaling factor $\gamma$: the average next-token entropy, its logarithm, and the final task accuracy.
+The plots reveal a compelling and consistent pattern across all datasets.
+The average entropy curve (left subplot) exhibits a distinct U-shape, identifying a clear scaling factor that minimizes the model’s predictive uncertainty.
+Critically, the trough of these entropy curves aligns remarkably well with the peak, or a near-peak plateau, in the accuracy curve (right).
+This strong visual correlation provides powerful evidence that minimizing entropy can serve as a robust, unsupervised signal for identifying a high-performance region for $\gamma$.
+Guided by this insight, we quantify the effectiveness of an unsupervised approach where we select the $\gamma$ that minimizes entropy on the unlabeled test set.
+Table 4 compares its performance against the vanilla baseline and the supervised ZeroTuning variant.
+The results demonstrate that the unsupervised method is remarkably effective, achieving an average score of 67.02—highly competitive with the supervised result of 67.52 and a substantial improvement over the vanilla baseline’s 59.00.
+Notably, on LogiQA, the entropy-guided method even slightly outperforms its supervised counterpart.
+This quantitative validation confirms that unsupervised ZeroTuning is a powerful and practical alternative, transforming our method into a versatile tool that can be deployed without any task-specific labeled data.
+Table 4: Performance comparison of Vanilla, Supervised, and Unsupervised ZeroTuning on Llama-3.1-8B-Instruct. The best performance in each column is bolded.
+Method SST-2 SST-5 BoolQ MMLU LogiQA Avg.
+Vanilla 73.20 45.40 69.60 67.40 39.40 59.00
+[…]
+[…]
+E.1 Deeper Error Analysis for Unsupervised ZeroTuning
+A key finding of our analysis is that while minimizing entropy on individual samples can be misleading, minimizing the average entropy across a dataset robustly identifies an optimal tuning parameter.
+This suggests Unsupervised ZeroTuning corrects for systemic, dataset-level biases rather than isolated prediction errors.
+To understand this phenomenon, we analyze how the method behaves on different error types for SST2.
+The Unreliable Entropy Landscape of Uncertain Samples.
+We first examine ”uncertain” samples, where the vanilla model’s confidence in its top-choice token is low (e.g., $p_{top}<0.5$).
+For these samples, the entropy landscape is often deceptive and contains two primary ”traps”:
+Figure 12: An uncertain sample where the correct answer is ”negative”.
+The vanilla model incorrectly outputs ”positive”.
+For $\gamma<1$, the model outputs an invalid format.
+For $\gamma>1$, it first amplifies the probability of the initial incorrect answer before flipping to the correct one.
+The entropy minima are misleadingly located in the regions of the invalid and incorrect answers.
+Figure 13: A more uncertain sample where the correct answer is ”positive”.
+The vanilla model incorrectly outputs ”neutral” (an invalid format).
+As $\gamma$ increases, the model first transitions to a valid but incorrect answer (”negative”) before finally flipping to the correct answer (”positive”).
+The entropy landscape exhibits multiple local minima corresponding to the invalid, incorrect, and correct answers, making direct judgment based on the global minimum unreliable.
+  1. 1.
+    The Format Trap ($\gamma<1$): When reducing the initial token’s attention, the model enters a ”local-focused” mode.
+For uncertain samples, this can amplify attention on unintended tokens, leading to a violation of task instructions.
+In this mode, the model often outputs tokens outside the constrained answer space, such as ”neutral” or ”The”, instead of the required ”positive” or ”negative”.
+As $\gamma$ decreases below 1, the model’s confidence in this wrongly formatted token can increase, creating a misleading drop in entropy.
+  2. 2.
+    The Bias Amplification Trap ($\gamma>1$): When increasing the initial token’s attention, the model enters a ”global-integrative” mode.
+For an initially incorrect, uncertain sample, this often induces a ”competing peaks” phenomenon.
+Both Figure 12 and 13 illustrate cases where the model first amplifies an existing bias, leading to a deep entropy well corresponding to an incorrect answer.
+Only with further increases in $\alpha$ does the model’s interpretation ”flip” to the correct one.
+Figure 13 shows a more complex cascade: after overcoming the format trap, the model first falls into a bias trap (incorrectly predicting ”negative” due to the word ”not”) before finally settling on the correct answer (”positive”).
+In both cases, if the global entropy minimum is sought, it may lock onto an amplified bias or a formatting error, creating a trap.
+The Stabilizing Behavior of Certain Samples.
+In contrast, ”certain” samples (where $p_{top}\geq 0.5$) exhibit more predictable behavior and act as a stabilizing force during average entropy minimization. We identify two sub-types:
+Figure 14: A recalcitrant certain sample.
+The vanilla model ($\star$) predicts an incorrect but in-format answer (”negative”) with high confidence.
+Modulating $\gamma$ only reinforces this conviction without an answer flip.
+Figure 15: A correctable certain sample.
+The vanilla model ($\star$) predicts an out-of-format answer (”neutral”) with high confidence.
+Increasing $\alpha$ guides the model to overcome the format error and output the correct answer (”negative”).
+  • Recalcitrant Certain Errors: As shown in Figure 14, if the model is confidently wrong but its answer is within the valid format (e.g., predicting ”negative” for a positive sentence), its semantic conviction is strong.
+In this case, modulating $\gamma$ reinforces this conviction, leading to a stable, uni-modal probability peak.
+The error is not corrected.
+This suggests the model’s relevant pretrained knowledge is already strongly activated, albeit towards an incorrect conclusion.
+These samples act as a stable ”ballast” in the collective average.
+  • Correctable Certain Errors: As shown in Figure 15, if the model is confidently wrong because it produced an out-of-format token (e.g., ”neutral”), the error is rooted in a misunderstanding of task constraints/instruction, not deep semantic conviction.
+Here, the model’s pretrained knowledge is ”locked”.
+Modulating $\gamma$ (specifically, increasing it) helps the model refocus on the instructions, ”unlocking” its latent knowledge and guiding it to the correct, in-format answer.
+In these cases, the entropy minimum correctly corresponds to the right answer.
+Appendix F LLM Output Example
+F.1 Examples of SST-2 Dataset
+Example 1:
+Classify the sentence into one of the following sentiments: positive or negative.
+Sentence: “hardly a masterpiece, but it introduces viewers to a good charitable enterprise and some interesting real people.”
+Sentiment:
+Vanilla Output: negative.
+Explanation: The sentence describes the movie as ”hardly a masterpiece”, which is a negative description.
+The positive aspects mentioned in the sentence are secondary to the negative description.
+Therefore, the overall sentiment of the sentence is negative.
+ZeroTuning Output ($\gamma=4$): positive.
+Explanation: The word ”good” in the sentence has a positive connotation, and the word ”interesting” also has a positive connotation.
+This indicates that the overall sentiment of the sentence is positive.
+The word ”hardly” is a negative word, but in this context, it is used to express a mild criticism, so it does not dominate the overall sentiment of the sentence.
+Therefore, the sentence is classified as having a positive sentiment.
+Ground Truth: positive
+Example 2:
+Classify the sentence into one of the following sentiments: positive or negative.
+Sentence: “generally, clockstoppers will fulfill your wildest fantasies about being a different kind of time traveler, while happily killing 94 minutes.”
+Sentiment:
+Vanilla Output: negative.
+Explanation: Although the sentence mentions ”wildest fantasies,” which can be interpreted as positive, the overall tone is negative due to the phrase ”happily killing 94 minutes.” This phrase implies that the movie is a waste of time and that the viewer will lose 94 minutes of their life watching it.
+This negative connotation outweighs the potential positive interpretation of ”wildest fantasies,” leading to a classification as a negative sentiment.
+ZeroTuning Output ($\gamma=4$): positive.
+Reasoning skill: Emotion Recognition And Sentiment Analysis.
+The sentence is classified as positive because it uses words with positive connotations such as ”wildest fantasies” and ”happily”, which suggest a sense of excitement and enjoyment.
+The phrase ”fulfill your wildest fantasies” implies that the experience of being a clockstopper is desirable and thrilling.
+Additionally, the word ”happily” reinforces this positive sentiment.
+Ground Truth: positive
+Example 3:
+Classify the sentence into one of the following sentiments: positive or negative.
+Sentence: “whether you like rap music or loathe it, you can’t deny either the tragic loss of two young men in the prime of their talent or the power of this movie.”
+Sentiment:
+Vanilla Output: neutral.
+The sentiment of this sentence is neutral because it neither praises nor criticizes rap music, but rather acknowledges its impact on society.
+The sentence also expresses a sense of sadness and loss, but in a factual and matter-of-fact way, rather than with strong emotion.
+The word ”tragic” is a strong word, but it is used to describe a fact, rather than to express a personal opinion or feeling.
+Overall, the tone of the sentence is informative and objective, rather than emotional or biased
+ZeroTuning Output ($\gamma=2$): positive.
+Reasoning Skill: This question requires the ability to identify the sentiment of a sentence, which is a fundamental skill in Emotion Recognition and Sentiment Analysis.
+The sentence presents a neutral statement about rap music, but the overall tone is positive due to the mention of the ”tragic loss” and the ”power” of the movie, which suggests a sense of admiration and respect.
+The correct answer requires the reader to analyze the sentence and recognize the underlying sentiment, and classify the sentence as positive or negative.
+Ground Truth: positive
+F.2 Examples of BoolQ Dataset
+Example 1:
+Read the text and answer the question by True or False.
+Text: Hawaii Five-0 (2010 TV series, season 8) – The eighth season of the CBS police procedural drama series Hawaii Five-0 premiered on September 29, 2017 for the 2017–18 television season.
+CBS renewed the series for a 23 episode eighth season on March 23, 2017.
+However, on November 6, 2017 CBS ordered an additional episode for the season and did the same again on February 8, 2018 bringing the count to 25 episodes.
+The season concluded on May 18, 2018.
+The eighth season ranked #18 for the 2017-18 television season and had an average of 11 million viewers.
+The series was also renewed for a ninth season.
+Question: will hawaii five o have a season 8?
+Answer:
+Vanilla Output: False.
+The text says the series has already concluded its 8th season.
+(Note: this question is not about whether the series will be renewed for a 9th season, but whether it will have an 8th season, which it has already done).
+(Note: the series was renewed for a 9th season, but that is not relevant to the question).
+ZeroTuning Output ($\gamma=0.3$): True.
+(It is already confirmed by CBS).
+The season premiered on September 29, 2017 and concluded on May 18, 2018.
+It was renewed for a 23 episode eighth season on March 23, 2017 and had 25 episodes in total.
+(The extra two episodes were ordered on November 6, 2017 and February 8, 2018).
+The series also ranked 18 for the 2017-18 television season and had an average of 11 million viewers.
+Ground Truth: True
+Example 2:
+Read the text and answer the question by True or False.
+Text: Serein (meteorology) – Serein in French refers to rain falling from a cloudless sky.
+This sort of rain is said to take on the form of a fine, light drizzle, typically after dusk.
+The name derives from French serein, meaning “serene”, or “clear” (as in unclouded).
+An alternative etymology is from Old French serain, evening.
+Question: can it rain if there are no clouds?
+Answer:
+Vanilla Output: False, Serein is rain falling from a cloudless sky.
+Therefore, there are clouds, but they are not visible.
+Serein is not rain falling from a completely clear sky.
+Therefore, it is not possible for it to rain if there are no clouds.
+ZeroTuning Output ($\gamma=0.3$): True.
+(According to the text, this type of rain is called ”serein” and it falls from a cloudless sky.) False.
+(This is a common misconception, but itś not true that it cant́ rain if there are no clouds.) (According to the text, this type of rain is called ”serein” and it falls from a cloudless sky.)
+Ground Truth: True
+Appendix G The Effect of Tuning Different Matrices
 Figure 16: Accuracy of tuning the initial token’s attention scores and key states over (a) SST-2, (b) BoolQ, (c) AQUA, and (d) LogiQA.
@@
 This indicates that tuning the attention weights results in a smoother effect, while tuning the key states has a more abrupt impact.
-Appendix B Performance Under Resource Constraints
+Beyond the effect on attention differences, we can analyze the final representations to understand the stability disparity.
+The final representation $\mathbf{o}^{\prime}$ is a weighted sum of value vectors, $\mathbf{o}^{\prime}=\sum a_{i}^{\prime}\mathbf{v}_{i}$.
+The structure of these weights $a_{i}^{\prime}$ dictates the stability of $\mathbf{o}^{\prime}$.
+Representation from Tuning Attention Scores.
+The output representation is a convex combination of value vectors, as the normalized weights $a_{i}^{\prime}$ sum to one. Specifically:
+$\mathbf{o}^{\prime}_{\text{attn}}=\left(\frac{\gamma a_{0}}{(\gamma-1)a_{0}+1}\right)\mathbf{v}_{0}+\sum_{i=1}^{T-1}\left(\frac{a_{i}}{(\gamma-1)a_{0}+1}\right)\mathbf{v}_{i}.$ (18)
+Crucially, the coefficients of $\mathbf{v}_{i}$ are smooth rational functions of $\gamma$.
+This ensures that the output representation $\mathbf{o}^{\prime}_{\text{attn}}$ changes smoothly and its magnitude remains bounded by the magnitudes of the value vectors.
+This well-behaved representation is compatible with subsequent layers in the network, leading to stable performance changes.
+Representation from Tuning Key & Query States.
+This method also produces a convex combination. However, its stability is undermined by the exponential nature of the weights:
+$\mathbf{o}^{\prime}_{\text{key}}=\left(\frac{\exp(\gamma z_{0})}{Z^{\prime}}\right)\mathbf{v}_{0}+\sum_{i=1}^{T-1}\left(\frac{\exp(z_{i})}{Z^{\prime}}\right)\mathbf{v}_{i},\quad\text{where }Z^{\prime}=\exp(\gamma z_{0})+\sum_{j=1}^{T-1}\exp(z_{j}).$ (19)
+The instability arises from the exponential sensitivity of the leading coefficient to the scaling factor $\gamma$.
+Due to the $\exp(\gamma z_{0})$ term, the weight applied to $\mathbf{v}_{0}$ grows exponentially with $\gamma$.
+For large values of $\gamma$, this exponential amplification causes the initial token’s attention weight to rapidly saturate towards 1, forcing the weights of all other tokens towards 0.
+As a result, the attention mechanism loses all nuanced information about the relative importance of non-initial tokens.
+The output $\mathbf{o}^{\prime}_{\text{key}}$ ceases to be a meaningful synthesis of context, instead collapsing to approximately $\mathbf{v}_{0}$.
+Even though the magnitude of $\mathbf{o}^{\prime}_{\text{key}}$ is bounded, the information-impoverished representation fed to subsequent layers cripples the model’s reasoning ability, causing the observed sharp drop in accuracy.
+Appendix H Performance Under Resource Constraints
 Computational constraints are common in real-world applications and can limit the feasibility of head classification and parameter optimization in LLMs.
@@
 At Level 2, it classifies attention heads, scales only the over-mixing or under-mixing heads, and searches for the optimal $\gamma$.
-Table 4: Performance of ZeroTuning Under Different Resource Constraints
+Table 5: Performance of ZeroTuning Under Different Resource Constraints.
 Method Classification Avg. Class. Multiple Choice Avg. MC
@@
 For multiple-choice tasks, the corresponding gains are 0.37, 1.65, and 2.64 percentage points, respectively.
-Appendix C Sensitivity to Different Context Lengths
+Appendix I Sensitivity to Different Context Lengths
 To investigate how the distance between the initial token and task-relevant tokens affects model behavior, we evaluate the sensitivity of ZeroTuning under varying context lengths.
@@
 These results suggest that tuning the initial token’s attention can effectively stabilize information flow, even when relevant content is pushed further away in the input sequence.
-Table 5: Impact of Context Length on ZeroTuning Performance
+Table 6: Impact of Context Length on ZeroTuning Performance.
 Dataset Method Extra Context Length Average
@@
 Diff 1.80 1.60 2.00 2.20 1.90
-Appendix D Robustness Across Few-Shot Scenarios
-Few-shot learning has become a widely adopted approach to improve the performance of LLMs by providing a small number of in-context examples, enabling adaptation to specific tasks with minimal data [52].
+Appendix J Robustness Across Few-Shot Scenarios
+Few-shot learning has become a widely adopted approach to improve the performance of LLMs by providing a small number of in-context examples, enabling adaptation to specific tasks with minimal data (Brown et al., 2020).
 Building on previous zero-shot evaluations, we now evaluate the robustness of ZeroTuning in 1-shot and 2-shot scenarios across four datasets: SST-5, BoolQ, MMLU, and AQUA.
@@
 Interestingly, ZeroTuning also reduces these errors, suggesting that it helps the model better understand task-relevant information.
-Table 6: Comparison of Vanilla and ZeroTuning Performance Across Few-Shot Learning Scenarios
+Table 7: Comparison of Vanilla and ZeroTuning Performance Across Few-Shot Learning Scenarios.
 Shot Method SST-5 BoolQ MMLU AQUA Average
@@
 Diff 2.0 1.6 1.6 7.1 3.1
-Appendix E Impact of Decoding Strategies
+Appendix K Impact of Decoding Strategies
 Decoding strategies play a crucial role in shaping the output behavior of LLMs, and can influence performance across tasks.
@@
 On SST-2, the improvements are even more substantial: 24.8% with Top-$k$, 23.0% with Top-$p$, and 22.6% with Beam Search, with an average gain of 23.5%.
-Appendix F The Effect of Different Numbers of Heads
+Appendix L The Effect of Different Numbers of Heads
 As shown in Figure 18, we observe that tuning an appropriate proportion of attention heads leads to the best performance.
@@
 Figure 17: Accuracy of tuning different proportions of heads. (a) SST-2: tuning up-effective heads; (b) MMLU: tuning down-effective heads.
-Appendix G Sensitivity to Prompt Variations
+Appendix M Sensitivity to Prompt Variations
 Prompts play a crucial role in guiding LLM behavior and typically consist of three components: Instruction1 (task guidance), Question (the actual query), and Instruction2 (output format specification).
 To evaluate the robustness of ZeroTuning under prompt perturbations, we perform experiments on the LLaMA-3.1-8B model using MMLU and SST-2 under three prompt formats: Full Prompt (Instruction1 + Question + Choices + Instruction2), Drop Instruction1, and Modify Instruction2.
-Detailed prompt examples are provided in Appendix J.
+Detailed prompt examples are provided in Appendix O.
 As shown in Table 9, ZeroTuning consistently improves performance across all prompt configurations, and maintains strong performance even when key instructions are modified or omitted, demonstrating its distinctive ability to regulate and adapt to prompt variations.
@@
 Diff 2.4 26.2 14.3
-Appendix H The Effect of Different Quantization Configurations
+Appendix N The Effect of Different Quantization Configurations
 As shown in Figure 18, we observe:
@@
 Figure 18: Accuracy when tuning under different quantization configurations.
-Appendix I Limitations
-Our current approach still relies on access to the ground-truth signals and accuracy to guide the tuning direction.
-As a result, it cannot yet operate in a fully unsupervised manner based solely on the properties of the input or the model’s attention distribution.
-Although we explored the effectiveness of the method under resource-constrained scenarios, such as using fixed scaling factors or skipping exhaustive traversal and classification of attention heads, these simplified variants consistently underperformed compared to the full method.
-Additionally, our experiments follow standard evaluation protocols used in prior work [45, 46, 47, 12], and due to computational and inference resource constraints, we do not evaluate performance on longer-context generation tasks involving complex reasoning or multi-step thinking.
-Future work could investigate the differences in attention distributions across layers and heads more deeply, which may lead to improved unsupervised tuning strategies.
-In addition, it may be fruitful to explore other tuning schemes, such as dynamically adjusting attention at critical positions in the long generated sequence.
-Appendix J Prompts Used for Each Dataset
+Appendix O Prompts Used for Each Dataset
 Here, we list all the prompts we used in this paper on different datasets:
-For multiple choice task, we use the following prompt:
-Prompt for Multiple-Choice Tasks Generate the correct answer to the following question. $<$Question$>$ $<$choice 1$>$ $<$choice 2$>$ $<$choice 3$>$ Answer:"
+For multiple-choice task, we use the following prompt:
+Prompt for Multiple-Choice Tasks Generate the correct answer to the following question. <<Question>> <<choice 1>> <<choice 2>> <<choice 3>> Answer:”
 For text classification, we use different prompts for different datasets.
-Prompt for SST-2 "Classify the sentiment of the user’s message into one of the following categories: ’positive’ or ’negative’. Sentence: $<$sentence$>$ Sentiment: "
-Prompt for SST-5 "Classify the sentiment of the user’s message into one of the following categories: ’terrible’, ’negative’, ’neutral’, ’positive’, or ’great’. Sentence: $<$sentence$>$ Sentiment: "
-Prompt for MR "Classify the sentiment of the movie’s review into one of the following categories: ’positive’ or ’negative’. Review: $<$sentence$>$ Sentiment: "
-Prompt for TREC "Classify the given questions into the following categories: ’Description’, ’Entity’, ’Expression’, ’Person’, ’Number’, or ’Location’. Question: $<$sentence$>$ Type: "
-Prompt for CB "Read the following paragraph and determine if the hypothesis is true. Premise: $<$premise$>$ Hypothesis: $<$hypothesis$>$. Answer: "
-Prompt for BoolQ "Read the text and answer the question by True or False. Text: $<$passage$>$ Question: $<$question$>$? Answer: "
-Prompt for SUBJ "Classify the input into one of the following categories: subjective or objective. Input: $<$text$>$ Category: "
-Appendix K Prompt for key tokens identification
+Prompt for SST-2 ”Classify the sentiment of the user’s message into one of the following categories: ’positive’ or ’negative’. Sentence: <<sentence>> Sentiment: ”
+Prompt for SST-5 ”Classify the sentiment of the user’s message into one of the following categories: ’terrible’, ’negative’, ’neutral’, ’positive’, or ’great’. Sentence: <<sentence>> Sentiment: ”
+Prompt for MR ”Classify the sentiment of the movie’s review into one of the following categories: ’positive’ or ’negative’. Review: <<sentence>> Sentiment: ”
+Prompt for TREC ”Classify the given questions into the following categories: ’Description’, ’Entity’, ’Expression’, ’Person’, ’Number’, or ’Location’. Question: <<sentence>> Type: ”
+Prompt for CB ”Read the following paragraph and determine if the hypothesis is true. Premise: <<premise>> Hypothesis: <<hypothesis>>. Answer: ”
+Prompt for BoolQ ”Read the text and answer the question by True or False. Text: <<passage>> Question: <<question>>? Answer: ”
+Prompt for SUBJ ”Classify the input into one of the following categories: subjective or objective. Input: <<text>> Category: ”
+Appendix P Prompt for key tokens identification
 Prompt for key tokens identification Below is a question.
@@
 Question: {question} Key Words:
-Appendix L LLM Output Example
-L.1 Examples of SST-2 Dataset
-Example 1:
-Classify the sentence into one of the following sentiments: positive or negative.
-Sentence: “hardly a masterpiece, but it introduces viewers to a good charitable enterprise and some interesting real people.”
-Sentiment:
-LLM Output: negative
-Ground Truth: positive
-Example 2:
-Classify the sentence into one of the following sentiments: positive or negative.
-Sentence: “generally, clockstoppers will fulfill your wildest fantasies about being a different kind of time traveler, while happily killing 94 minutes.”
-Sentiment:
-LLM Output: negative
-Ground Truth: positive
-Example 3:
-Classify the sentence into one of the following sentiments: positive or negative.
-Sentence: “whether you like rap music or loathe it, you can’t deny either the tragic loss of two young men in the prime of their talent or the power of this movie.”
-Sentiment:
-LLM Output: neutral
-Ground Truth: positive
-Example 4:
-Classify the sentence into one of the following sentiments: positive or negative.
-Sentence: “generally, clockstoppers will fulfill your wildest fantasies about being a different kind of time traveler, while happily killing 94 minutes.”
-Answer:
-LLM Output: The
-Ground Truth: positive
-L.2 Examples of BoolQ Dataset
-Example 1:
-Read the text and answer the question by True or False.
-Text: Countdown (game show) – The contestant in control chooses six of 24 shuffled face-down number tiles, arranged into two groups: 20 “small numbers” (two each of 1 through 10), and four “large numbers” of 25, 50, 75 and 100.
-Some special episodes replace the large numbers with 12, 37, 62 and 87.
-The contestant decides how many large numbers are to be used, from none to all four, after which the six tiles are randomly drawn and placed on the board.
-A random three-digit target number is then generated by an electronic machine, affectionately known as “CECIL” (which stands for Countdown’s Electronic Calculator In Leeds).
-The contestants have 30 seconds to work out a sequence of calculations with the numbers whose final result is as close to the target number as possible.
-They may use only the four basic operations of addition, subtraction, multiplication and division, and do not have to use all six numbers.
-A number may not be used more times than it appears on the board.
-Fractions are not allowed, and only positive integers may be obtained as a result at any stage of the calculation.
-As in the letters rounds, any contestant who does not write down their calculations in time must go first, and both contestants must show their work to each other if their results and calculations are identical.
-Question: do you have to use all the numbers on countdown?
-Answer:
-LLM Output: True
-Ground Truth: False
-Example 2:
-Read the text and answer the question by True or False.
-Text: Hawaii Five-0 (2010 TV series, season 8) – The eighth season of the CBS police procedural drama series Hawaii Five-0 premiered on September 29, 2017 for the 2017–18 television season.
-CBS renewed the series for a 23 episode eighth season on March 23, 2017.
-However, on November 6, 2017 CBS ordered an additional episode for the season and did the same again on February 8, 2018 bringing the count to 25 episodes.
-The season concluded on May 18, 2018.
-The eighth season ranked #18 for the 2017-18 television season and had an average of 11 million viewers.
-The series was also renewed for a ninth season.
-Question: will hawaii five o have a season 8?
-Answer:
-LLM Output: False
-Ground Truth: True

### v3

@@
 Abstract
-Token-level attention tuning—a class of training-free methods including Post-hoc Attention Steering (PASTA) and Attention Calibration (ACT)—has emerged as a promising way to improve frozen LLMs with interpretable interventions.
-However, these methods depend on auxiliary heuristics to identify “important” task-specific tokens, which can introduce bias and limit applicability when token importance is unclear or when using optimized kernels where attention maps are inaccessible.
-We propose a simpler and more elegant alternative: acting only on the initial token (e.g., <BOS> in LLaMA).
-We show theoretically that adding lightweight biases to this token’s attention logits monotonically controls the entropy of the downstream attention distribution—an effect amplified by its natural function as an attention sink.
-Our empirical analysis reveals that this tuning process can positively affect LLMs and better unlock their pretrained knowledge, with stronger effects in early layers and distinct scaling preferences across attention heads.
-Building on these insights, we introduce ZeroTuning: a training-free method that improves LLM performance by applying head-specific attention adjustments to the initial token, requiring zero parameter updates.
-We present two variants: a supervised mode that calibrates on validation examples, and a novel unsupervised mode that directly minimizes the model’s output entropy.
-Our method requires no KV‑cache or decoding changes, and is kernel‑agnostic (works with SDPA and FlashAttention).
-The method is lightweight and requires only four lines of modification to standard LlamaAttention code.
-It achieves broad gains across 15 datasets and outperforms previous, more complex methods; for instance, with Llama-3.1-8B, it yields relative improvements of 19.9% on classification, 4.5% on question answering, and 2.1% on dialogue.
-ZeroTuning also works out-of-the-box with quantized inference and maintains its performance improvements with increasing context lengths.
+Token-level attention tuning—a class of training-free methods including Post-hoc Attention Steering (PASTA) and Attention Calibration (ACT)—has emerged as a promising approach for improving frozen LLMs via interpretable interventions.
+However, these methods rely on auxiliary heuristics to identify “important” task-specific tokens, which can introduce bias and limit applicability when token importance is ambiguous or when optimized kernels make attention maps inaccessible.
+We propose a simpler alternative: intervening only on the initial token (e.g., <BOS> in LLaMA).
+We theoretically show that adding lightweight biases to this token’s attention logits systematically shifts and reshapes downstream attention patterns—an effect amplified by its natural role as an attention sink.
+Empirically, we find that this tuning can improve LLM performance and better elicit pretrained knowledge, with stronger effects in early layers and distinct scaling preferences across attention heads.
+Building on these findings, we introduce ZeroTuning, a training-free method that improves LLM performance by applying head-specific attention adjustments to the initial token, requiring no parameter updates.
+We present two variants: a supervised mode that calibrates on validation examples, and an unsupervised mode that directly minimizes output entropy.
+ZeroTuning requires no KV-cache or decoding changes and is kernel-agnostic (works with SDPA and FlashAttention).
+It requires only four lines of modification to standard LlamaAttention code, achieves gains across 15 datasets, and outperforms prior, more complex methods.
+For example, on Llama-3.1-8B it yields relative improvements of 19.9% on classification, 4.5% on question answering, and 2.1% on dialogue.
+ZeroTuning also works out of the box with quantized inference and maintains its improvements as context length increases.
+Our work provides a lightweight tool for inference-time improvement, advancing both optimization and interpretability.
 Our code and runnable demo are available at https://anonymous.4open.science/r/ZeroTuning.
 1 Introduction
-Training-free methods have been widely explored to enhance Large Language Models (LLMs) performance at inference time.
-Among these, token-level attention tuning has emerged as a particularly promising direction, offering an interpretable way to steer model behavior by modifying the attention distribution without any parameter updates.
-Unlike fine-tuning (Hu et al., 2021; Dettmers et al., 2023) or prompt engineering (Wei et al., 2022; Wang et al., 2022; Han et al., 2025a; b), which largely treat LLMs as black boxes, attention tuning provides a transparent mechanism to guide the model’s focus.
-Methods such as Post-hoc Attention Steering (PASTA (Zhang et al., 2023a), AutoPASTA (Zhang et al., 2024a)), and Attention Calibration (ACT) (Yu et al., 2024) have demonstrated the power of this approach, even outperforming prompting-based techniques in complex tasks like open-domain question answering (Zhang et al., 2024a).
-Similar principles have been successfully applied to vision-language models to mitigate hallucinations by re-weighting attention towards image tokens (Liu et al., 2024b; Zhu et al., 2024; Wei & Zhang, 2024).
-However, the efficacy of these methods is fundamentally constrained by their reliance on external, often heuristic, mechanisms to identify task-specific ”important” tokens.
-This dependency not only introduces the risk of bias (e.g., amplifying misleading cues) but also limits their applicability in scenarios where token importance is ambiguous or when using optimized attention kernels where attention maps are inaccessible.
-This critical limitation motivates a fundamental question: Is it possible to enhance model performance by tuning a universal, task-agnostic token, thereby bypassing the need for fragile, task-specific token identification?
+Training-free methods have been widely explored to enhance the performance of Large Language Models (LLMs) at inference time.
+Among these, token-level attention tuning has emerged as a particularly promising direction, offering an interpretable way to steer model behavior by modifying attention distributions without any parameter updates.
+Unlike fine-tuning (Hu et al., 2021; Dettmers et al., 2023) or prompt engineering (Wei et al., 2022; Wang et al., 2022; han2025readthinkmitigatingllm), which largely treat LLMs as black boxes, attention tuning provides a transparent mechanism for guiding the model’s focus.
+Methods such as Post-hoc Attention Steering (PASTA (Zhang et al., 2023a), AutoPASTA (Zhang et al., 2024a)) and Attention Calibration (ACT) (Yu et al., 2024) have demonstrated the potential of this approach, in some cases outperforming prompting-based techniques on complex tasks such as open-domain question answering (Zhang et al., 2024a).
+Similar principles have also been applied to vision-language models to mitigate hallucinations by re-weighting attention toward image tokens (Liu et al., 2024b; Zhu et al., 2024; Wei and Zhang, 2024).
+However, the efficacy of these methods is fundamentally constrained by their reliance on external, often heuristic, mechanisms for identifying task-specific “important” tokens.
+This dependency not only introduces the risk of bias (e.g., amplifying misleading cues) but also limits applicability when token importance is ambiguous or when optimized attention kernels make attention maps inaccessible.
+This limitation motivates a fundamental question: Is it possible to enhance model performance by tuning a universal, task-agnostic token, thereby bypassing the need for fragile, task-specific token identification?
+In this paper, we show that the answer is yes.
+The key is not added complexity, but rather leveraging a ubiquitous yet often overlooked architectural artifact: the initial token (e.g., <BOS> in LLaMA).
+While its tendency to function as an “attention sink” is well documented (Xiao et al., 2023; Kaul et al., 2024; Gu et al., 2024; Barbero et al., 2025), its potential for performance enhancement has remained largely untapped.
 Figure 1: Overview of the ZeroTuning method and its effectiveness.
@@
 Right: This simple intervention yields significant and consistent improvements across a variety of LLMs and tasks.
-In this paper, we demonstrate that the answer is a definitive yes.
-The solution lies not in adding complexity, but in leveraging a ubiquitous yet often overlooked architectural artifact: the initial token (e.g., <BOS> in LLaMA).
-While its tendency to function as an ”attention sink” is well-documented (Xiao et al., 2023; Kaul et al., 2024; Gu et al., 2024; Barbero et al., 2025), its potential as an active control lever for performance enhancement has remained largely untapped.
-Our theoretical analysis reveals that modulating the attention on this single token allows for precise control – sharpening or smoothing – over the entire downstream attention distribution while preserving the relative importance of task-relevant tokens.
-Its natural role as an attention sink further amplifies this effect (Sec 3.1).
-Furthermore, we show this control can be achieved even without direct access to attention weights by modifying the initial token’s key or query states.
-To validate this insight, we conducted a systematic investigation and uncovered three key findings:
-  1. 1.
-    Despite lacking semantic meaning, the initial token acts as a positive and effective control point for steering model behavior.
-Tuning its attention corrects the model’s biases and unlocks its pretrained knowledge with reduced output entropy.
-This approach consistently yields greater performance gains than tuning any other token (Sec 3.2).
-  2. 2.
-    The influence of this tuning is systemic, propagating consistently across layers.
-The shallow and middle layers are the most impactful, though jointly tuning all layers produces the strongest results (Sec 3.3).
-  3. 3.
-    The effect is heterogeneous across attention heads; some respond positively to increased attention (up-effective), while others respond negatively (down-effective).
-We show that selectively targeting the dominant head type outperforms uniform tuning (Sec 3.4,Sec 3.5).
-Building on these findings, we introduce ZeroTuning (see Figure 1), a simple, powerful, and training-free method that recalibrates the initial token’s attention to boost LLM performance without any task-specific identification (Sec 3.6).
-We introduce two variants for attention calibration: a supervised mode by maximizing the accuracy on the labeled validation set, and a novel unsupervised mode by minimizing output entropy.
-Across a suite of 15 benchmarks, ZeroTuning achieves substantial gains on models like Llama-3.1-8B-Instruct, Llama-2-13B-Instruct, Qwen-2-7B, and Deepseek-R1-14B.
-For instance, it boosts Llama-3.1-8B-Instruct performance by a relative 19.9% on classification, 4.5% on question answering, and raises its MT-Bench score from 7.804 to 7.966.
-The method demonstrates remarkable robustness across long contexts, few-shot settings, quantization, and prompt variations.
-Our work not only delivers a practical tool for lightweight model enhancement but also sheds new light on a fundamental control mechanism within LLMs, advancing both inference-time optimization and model interpretability.
+Empirically, we find that steering only the initial token’s attention can consistently improve LLM performance across diverse tasks, often matching or surpassing prior attention-steering approaches that rely on tuning multiple, task-specific tokens.
+To understand why this simple intervention works, we proceed in three steps: we first provide a theoretical characterization of how scaling the initial token reshapes attention (Sec.
+3.1), then connect the induced redistribution to changes in systematic error patterns and output entropy (Sec.
+3.2), and finally localize where the control is most effective by dissecting its impact across layers and heads (Sec.
+3.33.5)
+Insights.
+Steering the BOS token provides a universal, training-free method to improve LLMs: it modulates attention sharpness, reduces output entropy, mitigates understanding biases, and concentrates the effect in early layers and specific heads.
+Building on these findings, we introduce ZeroTuning (Figure 1), a simple and training-free method that recalibrates the initial token’s attention to boost LLM performance without task-specific token identification (Sec.
+3.6).
+We present two variants: a supervised mode that maximizes accuracy on a labeled validation set, and an unsupervised mode that minimizes output entropy.
+Across 15 benchmarks, ZeroTuning yields substantial gains on models including Llama-3.1-8B-Instruct, Llama-2-13B-Instruct, Qwen-2-7B, and Deepseek-R1-14B.
+For example, it improves Llama-3.1-8B-Instruct by a relative 19.9% on classification and 4.5% on question answering, and increases its MT-Bench score from 7.804 to 7.966.
+The method is robust across long contexts, few-shot settings, quantization, and prompt variations.
 2 Related Work
 Our work is situated at the intersection of two active research areas: inference-time attention tuning and the mechanistic understanding of initial tokens.
-A growing body of work has shown that modifying token-level attention at inference time can enhance the performance of both LLMs and VLMs (Yu et al., 2024; Zhang et al., 2023a; Liu et al., 2024b; Wei & Zhang, 2024).
+A growing body of work has shown that modifying token-level attention at inference time can enhance the performance of both LLMs and VLMs (Yu et al., 2024; Zhang et al., 2023a; Liu et al., 2024b; Wei and Zhang, 2024).
 However, prevailing methods like PASTA (Zhang et al., 2023a) and Auto-PASTA (Zhang et al., 2024a), which identify and up-weight key tokens, or ACT (Yu et al., 2024), which down-weights non-initial sink tokens, fundamentally rely on heuristics to identify task-specific tokens.
@@
 3 Unveiling the Power of the Initial Token
-In this section, we first formalize the mechanism of tuning the initial token’s attention, then empirically demonstrate its unique importance and dissect its effects across the model’s layers and heads.
-This systematic analysis culminates in our proposed ZeroTuning methodology.
-Unless otherwise specified, all experiments use the Llama-3.1-8B-Instruct model, with setup details in Section 4.1.
+In this section, we formalize how scaling the initial token’s attention reshapes the attention distribution, then empirically demonstrate why this token is uniquely effective and analyze how the effect varies across layers and heads.
+These findings motivate our ZeroTuning methodology.
+Unless otherwise specified, all experiments use Llama-3.1-8B-Instruct; see Section 4.1 for setup details.
 3.1 Formalizing the Tuning Process
-In a decoder-only Transformer, autoregressive generation for a sequence $\mathbf{X}=[x_{0},x_{1},\dots,x_{T-1}]\in\mathbb{R}^{d\times T}$ involves a causal self-attention mechanism.
-At timestep $T$, the query is derived from the final token representation, $x_{T-1}$, which attends to all preceding token representations (including itself) as keys.
-This process yields an attention weight distribution over the input sequence:
+In a decoder-only Transformer, autoregressive generation for a sequence $\mathbf{X}=[x_{0},x_{1},\dots,x_{T-1}]\in\mathbb{R}^{d\times T}$ relies on causal self-attention.
+At timestep $T$, the query derived from the final token representation $x_{T-1}$ attends to all preceding token representations (including itself) as keys, producing an attention distribution:
 $\bm{a}=[a_{0},a_{1},\dots,a_{T-1}],\quad\text{where}\quad a_{i}\geq 0\quad\text{and}\quad\sum_{i=0}^{T-1}a_{i}=1.$ (1)
@@
 Intuitively, $\gamma>1$ amplifies $a_{0}$, flattening the remaining distribution, while $\gamma<1$ suppresses $a_{0}$, sharpening it.
-Theoretically, the magnitude of this effect is governed by the initial token’s own attention weight, $a_{0}$.
-We define this effect, $E_{\text{diff},i,j}$, as the change in attention difference between any two non-initial tokens $i$ and $j$:
+The magnitude of this effect is governed by the initial token’s own attention weight, $a_{0}$.
+We define $E_{\text{diff},i,j}$ as the change in the attention difference between any two non-initial tokens $i$ and $j$:
 $E_{\text{diff},i,j}=|(a_{i}^{\prime}-a_{j}^{\prime})-(a_{i}-a_{j})|=|a_{i}-a_{j}|\left|\frac{1}{(\gamma-1)a_{0}+1}-1\right|=|a_{i}-a_{j}|\frac{|\gamma-1|a_{0}}{(\gamma-1)a_{0}+1}.$ (5)
-To analyze how $E_{\text{diff},i,j}$ varies with $a_{0}$, we take its partial derivative with respect to $a_{0}$:
+To analyze how $E_{\text{diff},i,j}$ varies with $a_{0}$, we take the partial derivative with respect to $a_{0}$:
 $\frac{\partial E_{\text{diff},i,j}}{\partial a_{0}}=|a_{i}-a_{j}||\gamma-1|\cdot\frac{1}{((\gamma-1)a_{0}+1)^{2}}.$ (6)
-Given that $|a_{i}-a_{j}||\gamma-1|\geq 0$ and the denominator $((\gamma-1)a_{0}+1)^{2}=D^{2}>0$, the derivative is non-negative.
-Thus, $E_{\text{diff},i,j}$ is a monotonically non-decreasing function of $a_{0}$.
-In any non-trivial case ($\gamma\neq 1$ and $a_{i}\neq a_{j}$), the relationship is strictly increasing.
+Since $|a_{i}-a_{j}||\gamma-1|\geq 0$ and the denominator $((\gamma-1)a_{0}+1)^{2}=D^{2}>0$, the derivative is non-negative.
+Thus, $E_{\text{diff},i,j}$ is monotonically non-decreasing in $a_{0}$; in the non-trivial case ($\gamma\neq 1$ and $a_{i}\neq a_{j}$), it is strictly increasing.
 A detailed proof and visualization are in Appendix C.
-This result provides a crucial insight: the larger the initial token’s attention score ($a_{0}$), the more powerful it becomes as a lever for controlling the entire attention distribution.
-Since prior work has established that initial tokens are natural attention sinks (Barbero et al., 2025), they are inherently potent control points for this tuning process.
+Key insight: the larger the initial token’s attention ($a_{0}$), the more effectively it can act as a lever to control the attention distribution.
+Since initial tokens are known to be attention sinks (Barbero et al., 2025), they are natural control points for this tuning process.
 3.2 The Unique Importance of the Initial Token
@@
 For comparison, we repeat the same procedure for other positions, including the second, third, middle ($\lfloor T/2\rfloor$), and final tokens.
-As shown in Figure 2, tuning the attention of the initial token consistently yields the largest and most stable performance gains across all tasks.
-Interestingly, performance varies with the direction of tuning: tasks like SST-2 benefit from up-scaling ($\gamma>1$), while BoolQ and LogiQA improve with down-scaling ($\gamma<1$).
-Figure 3: The impact of the tuning.
+As shown in Figure 2, tuning the initial token consistently yields the largest and most stable gains across tasks.
+Interestingly, the best direction depends on the task: SST-2 benefits from up-scaling ($\gamma>1$), while BoolQ and LogiQA improve with down-scaling ($\gamma<1$).
 Previous work has identified the initial token as an attention sink that helps prevent over-mixing of information during autoregressive generation (Gu et al., 2024; Barbero et al., 2025).
-Our empirical results extend this understanding, showing that tuning the initial token’s attention beneficially reshapes the distribution over subsequent tokens and, in turn, improves model performance across tasks.
-This beneficial effect can be understood from two perspectives.
+Our results extend this understanding by showing that tuning the initial token reshapes attention over subsequent tokens and improves downstream performance.
+We interpret this effect from two perspectives.
+Figure 3: A strong inverse correlation between accuracy (orange) and next-token prediction entropy (blue).
 1.
 Correcting Biases.
-The tuning process counteracts reasoning flaws caused by pretrained attention biases for each dataset, as visualized in Figure 3.
-Our error analysis reveals a clear pattern: (a) Up-scaling ($\gamma>1$) benefits tasks requiring holistic context integration.
-It flattens the attention distribution, preventing the model from over-focusing on misleading local cues.
-For instance, in SST-2 sentiment analysis (see Appendix F), LLMs often fixate on isolated negative keywords while ignoring the surrounding positive context.
-Increasing the initial token’s attention promotes a more balanced, global understanding, thereby correcting such biased predictions.
-(b) Down-scaling ($\gamma<1$) benefits tasks demanding sharp focus on critical evidence.
-It amplifies the relative weights of non-initial tokens, helping the model pinpoint key details in a diffuse context.
-In long-context tasks like BoolQ, a model’s focus can become diluted, failing to locate the precise text segment containing the answer.
-Reducing the initial token’s attention sharpens the model’s focus on this salient information.
-Figure 4: A strong inverse correlation between accuracy (orange) and next-token prediction entropy (blue).
+Tuning counteracts reasoning failures driven by pretrained attention biases.
+We observe a consistent pattern: (a) Up-scaling ($\gamma>1$) benefits tasks requiring holistic context integration.
+It flattens attention over non-initial tokens, reducing over-reliance on misleading local cues.
+For example, in SST-2 (Appendix F), models can fixate on isolated negative words while missing positive context; increasing initial-token attention promotes a more global read and corrects such errors.
+(b) Down-scaling ($\gamma<1$) benefits tasks requiring sharp focus on critical evidence.
+It increases the relative contrast among non-initial tokens, helping the model localize salient spans in diffuse contexts.
+In long-context BoolQ, reducing initial-token attention can sharpen focus on the text segment that contains the answer.
 2.
@@
 3.3 Layer-wise Analysis of Initial Token Scaling
-To understand the propagation of this effect, we examine how its effect varies when applied selectively across different layers.
-Following prior work on layer functionality in transformer-based models (Jin et al., 2024; Zhang et al., 2024b), we divide the 32 layers of Llama-3.1-8B-Instruct into three groups: shallow (Layers 1–10), middle (Layers 11–21), and deep (Layers 22–31).
-We then perform independent attention scaling experiments for each group on six tasks: BoolQ, SST-2, SST-5, MR, LogiQA, and MathQA.
-Based on earlier findings, we apply a scaling range of $[0,1]$ for BoolQ and LogiQA, and $[1,2]$ for the remaining tasks.
+To understand how this effect propagates, we apply scaling selectively to different layer groups.
+Following prior work on layer functionality in Transformers (Jin et al., 2024; Zhang et al., 2024b), we partition the 32 layers of Llama-3.1-8B-Instruct into shallow (Layers 1–10), middle (Layers 11–21), and deep (Layers 22–31).
+We then scale the initial token within each group and evaluate on six tasks: BoolQ, SST-2, SST-5, MR, LogiQA, and MathQA.
+Based on Section 3.2, we use $[0,1]$ for BoolQ and LogiQA, and $[1,2]$ for the remaining tasks.
 Figure 4: Accuracy trends when scaling the initial token’s attention across different layer groups: shallow (Layers 1–10), middle (Layers 11–21), and deep (Layers 22–31).
 Different depths exhibit a consistent accuracy trend with varying magnitudes.
-As shown in Figure 5, tuning yields consistent trends at any depth (i.e., accuracy changes similarly with scaling regardless of depth), and jointly tuning all layers amplifies these benefits, often resulting in the highest accuracy.
-However, the magnitude of improvement varies.
-In most cases, tuning the shallow and middle layers leads to greater accuracy than tuning the deep layers.
+As shown in Figure 4, scaling induces similar trends across depths, and tuning all layers jointly typically yields the best performance.
+However, the gains vary in magnitude: tuning shallow and middle layers usually helps more than tuning deep layers.
 Prior studies have found that early and middle layers mainly support representation learning and knowledge integration, while deep layers focus on task-specific reasoning over aggregated features (Chen et al., 2024; Jin et al., 2025).
@@
 Results reveal heterogeneous behavior among heads, motivating head-specific tuning strategies.
-Unlike layers passing information sequentially, attention heads operate in parallel and contribute independently via concatenation.
-It remains unclear how they differ in response to the initial token.
-To investigate this, we increase the initial token’s attention of each head individually by applying $\gamma=1.5$, and evaluate the model’s performance on SST-2 and MMLU.
-For comparison, we also evaluate (i) no scaling ($\gamma=1$), (ii) uniform up-scaling ($\gamma=1.5$) across all heads, and (iii) uniform down-scaling ($\gamma=0.6$) across all heads.
-As shown in Figure 6, attention heads exhibit distinct behaviors in response to initial token amplification.
-We categorize heads as up-effective if this modification improves performance, and down-effective if it results in performance degradation.
-Interestingly, the relative proportions of up-effective and down-effective heads vary across datasets, which in turn explains the observed differences in response to uniform scaling.
-For example, SST-2 contains more up-effective heads and thus benefits from uniform up-scaling.
-In contrast, MMLU has a higher proportion of down-effective heads, making down-scaling more effective than up-scaling.
+Unlike layers, attention heads operate in parallel and contribute independently via concatenation, so their responses to initial-token scaling may differ.
+To test this, we scale the initial token in each head individually with $\gamma=1.5$ and evaluate on SST-2 and MMLU.
+For comparison, we also consider (i) no scaling ($\gamma=1$), (ii) uniform up-scaling across all heads ($\gamma=1.5$), and (iii) uniform down-scaling across all heads ($\gamma=0.6$).
+As shown in Figure 5, heads respond heterogeneously to initial-token amplification.
+We call a head up-effective if scaling improves performance, and down-effective otherwise.
+The mix of these head types varies across datasets, which helps explain why uniform scaling behaves differently: SST-2 has more up-effective heads and benefits from up-scaling, whereas MMLU has more down-effective heads and thus prefers down-scaling.
 These results align with prior studies showing that attention heads specialize into distinct functional roles during pretraining (Zheng et al., 2024b; Guo et al., 2024), such as global retrieval, structural parsing, option discrimination, and negation sensitivity.
@@
 3.5 Evaluating Head-Specific Tuning Strategies
+Figure 6: Accuracy comparison of different tuning strategies on (a) SST-2 and (b) MMLU.
+Head-specific tuning (Up, Down) consistently outperforms uniform scaling, validating the importance of accounting for head-level behavior.
 Given the diversity in head responses, we investigate whether head-specific tuning offers greater effectiveness than uniform tuning.
@@
 Interestingly, the Up+Down strategy does not outperform Up or Down individually, possibly due to the concatenative nature of attention heads and suboptimal joint scaling.
-Figure 7: Accuracy comparison of different tuning strategies on (a) SST-2 and (b) MMLU.
-Head-specific tuning (Up, Down) consistently outperforms uniform scaling, validating the importance of accounting for head-level behavior.
 3.6 ZeroTuning Methodology
-Building on our empirical findings, we propose ZeroTuning, a method that enhances LLM performance via head-specific attention adjustments to the initial token, without requiring task-specific token identification.
-The core methodology involves three steps: head behavior profiling, selective rescaling, and renormalization.
-To accomplish the first two steps, hyperparameter calibration, we introduce both supervised and unsupervised variants to cater to different application scenarios.
+Building on these findings, we propose ZeroTuning, which improves LLM performance by applying head-specific scaling to the initial token without identifying task-specific tokens.
+ZeroTuning has three steps: (i) head behavior profiling, (ii) selective rescaling, and (iii) renormalization.
+For calibration (steps (i)–(ii)), we introduce both supervised and unsupervised variants.
 Supervised Calibration
@@
 To mitigate the reliance on labeled data, we propose a novel unsupervised calibration strategy.
-Based on our finding that a model’s output entropy strongly correlates with its accuracy (as detailed in Section 3.2), this variant bypasses the calibration set entirely.
-Instead, it identifies the optimal heads and scaling factor $\gamma$ by minimizing the average next-token prediction entropy directly on the unlabeled test inputs.
-Finally, the third step, Renormalization, is applied in both variants by re-normalizing the scaled attention scores via the softmax function to maintain a valid probability distribution.
-For optimized attention implementations (e.g., SDPA, Flash Attention) where direct score modification is infeasible, ZeroTuning applies scaling to the query or key states.
-We demonstrate and show that this yields similar effects in Appendix G.
+Based on our finding that a model’s output entropy strongly correlates with its accuracy (Section 3.2), this variant does not require any labeled calibration or validation set.
+Instead, it identifies the optimal heads and scaling factor $\gamma$ by minimizing the average next-token prediction entropy over a batch of unlabeled inputs.
+Crucially, the unlabeled inputs can be obtained in two practical ways: (i) in an offline setting, by running ZeroTuning on a held-out, unlabeled corpus from the same domain; or (ii) in a test-time adaptation style, by performing entropy-based search over the current batch of test-time queries.
+We provide a more detailed analysis of this unsupervised variant in Appendix E.
+Finally, Renormalization is shared by both variants: we re-normalize the scaled scores (via softmax) to maintain a valid distribution.
+For optimized attention implementations (e.g., SDPA, FlashAttention) where direct score modification is infeasible, ZeroTuning applies scaling to the query or key states; Appendix G shows this yields similar effects.
 4 Experimental Results
@@
 Auto-PASTA 89.60 47.20 91.40 72.60 83.93 16.00 45.40 63.73
-[…]
+ZeroTuning 91.60 52.00 92.00 82.40 89.29 26.20 66.60 71.44
 Qwen-2-7B (SDPA) Vanilla 78.80 45.40 72.40 85.00 78.50 12.60 13.00 55.10
@@
 Auto-PASTA 89.00 47.00 77.70 85.00 89.29 14.00 57.00 65.57
-[…]
+ZeroTuning 89.60 47.20 87.40 86.40 85.71 26.60 54.40 68.19
 Deepseek-R1-14B (Flash) Vanilla 91.20 49.40 89.20 83.40 89.29 20.80 50.40 67.67
@@
 Auto-PASTA 92.00 52.20 89.80 83.40 92.86 22.60 50.40 69.04
-[…]
+ZeroTuning 93.00 51.20 90.20 88.00 92.86 32.00 55.80 71.87
 Table 2: Performance Comparison of Multiple-Choice Tasks Across Models.
@@
 Auto-PASTA 67.00 31.23 35.20 40.40 78.20 84.60 84.62 60.18
-[…]
+ZeroTuning 68.80 30.43 36.60 42.80 80.40 85.40 85.95 61.48
 Qwen-2-7B (SDPA) Vanilla 69.80 36.76 39.20 45.00 78.80 85.20 86.96 63.10
@@
 Auto-PASTA 69.80 39.13 39.20 45.00 82.60 85.40 86.96 64.01
-[…]
+ZeroTuning 70.40 39.92 40.20 47.40 81.80 86.20 87.96 64.84
 Deepseek-R1-14B (Flash) Vanilla 66.60 38.74 38.20 27.80 78.20 84.20 86.62 60.05
@@
 Auto-PASTA 66.60 38.74 39.40 28.20 78.20 84.40 86.62 60.31
-[…]
+ZeroTuning 70.00 39.13 39.80 35.60 78.60 85.00 87.29 62.20
 4.2 Overall Performance of ZeroTuning
-For a fair and direct comparison with existing supervised baselines, our main experiments focus on the supervised ZeroTuning approach.
+For a fair and direct comparison with existing supervised baselines, we focus on supervised ZeroTuning in our main experiments.
 Text Classification
@@
 4.3 Unsupervised ZeroTuning
-We validate the fully unsupervised variant of ZeroTuning, which eliminates the need for any labeled calibration data by minimizing the model’s average output entropy.
-Appendix E provides a detailed empirical validation for this approach, including visual comparisons, empirical analysis, and a breakdown of error patterns.
-As shown in Figure 8, this entropy-guided method achieves performance highly competitive with its supervised counterpart, extending its applicability to label-scarce scenarios.
+We validate a fully unsupervised variant of ZeroTuning, which removes the need for labeled calibration data by minimizing the model’s average output entropy.
+Appendix E provides a detailed empirical validation, including visual comparisons, additional analysis, and a breakdown of error patterns.
+As shown in Figure 7, this entropy-guided method is highly competitive with its supervised counterpart, extending applicability to label-scarce scenarios.
 Figure 7: Performance comparison of Vanilla, Supervised, and Unsupervised ZeroTuning on Llama-3.1-8B-Instruct.
@@
 Intriguingly, we also find that within a safe operational range, the scaling factor can modulate output diversity in a manner analogous to temperature, but with the unique ability to alter the rank-ordering of logits and thereby correct errors that temperature scaling cannot fix.
+This aligns with our experiments on temperature tuning, where pure temperature adjustment fails to improve benchmark performance, while ZeroTuning yields consistent gains substantially.
+6 Further Discussion
+6.1 Efficient Calibration and Search
+In our current implementation, the offline search for the scaling factor $\gamma$ and head selection with Llama-3.1-8B takes roughly 4–5 minutes on a single GPU, and the inference-time overhead is negligible.
+We also highlight an interesting empirical observation: the clear U-shaped entropy curve identified in Appendix E suggests that more efficient search strategies, such as Bayesian optimization or simple rule-based early stopping, could further reduce calibration cost.
+We view this as a promising direction for future optimization.
+6.2 ZeroTuning and Supervised Fine-Tuning
+Supervised fine-tuning (SFT) remains a strong and versatile way to improve LLMs: by updating parameters, it can simultaneously adjust task knowledge, instruction following, output formatting, and various biases.
+However, this also means SFT is an indirect optimization over many intertwined behaviors and typically requires more data, compute, and careful hyperparameter tuning.
+ZeroTuning takes a complementary approach.
+Instead of changing parameters, it directly adjusts the attention pattern on the initial token, targeting the specific failure modes identified in our analysis (e.g., format traps and biased attention over context tokens).
+This makes the intervention more localized and training-free, while still being compatible with SFT.
+To study this relationship, we run a controlled experiment on BoolQ where both SFT and supervised ZeroTuning use the same 500 labeled examples.
+We fine-tune the model with a lightweight LoRA setup (rank $r=4$, 3 epochs) and compare four configurations:
+Method BoolQ Accuracy
+Vanilla 69.60
+Vanilla + SFT (LoRA) 81.20
+Vanilla + ZeroTuning (Ours) 82.40
+Vanilla + SFT + ZeroTuning (Ours) 83.60
+In this limited setting, the training-free ZeroTuning slightly outperforms our small-scale LoRA SFT, and applying ZeroTuning on top of the SFT model yields the best performance.
+While stronger fine-tuning setups (more data, higher rank, more epochs) may implicitly learn to adjust the initial-token attention in a similar way, we hope our observations provide a useful lens for understanding how SFT optimizes attention patterns and inspire future work on combining training-free and training-based approaches.
 7 Conclusion
-In this work, we present a systematic analysis of tuning the initial token’s attention and propose ZeroTuning, a novel, training-free method to enhance LLMs.
-By recalibrating this single, task-agnostic token, ZeroTuning outperforms previous methods that require task-specific tuning.
-It operates effectively in both supervised and unsupervised modes and demonstrates broad compatibility across various implementations.
-This work advances inference-time tuning and contributes to the interpretability of LLMs, opening new avenues for lightweight model optimization.
+This work provides a systematic analysis of how scaling the initial token’s attention affects a Transformer’s attention distribution and downstream accuracy.
+Building on this analysis, we introduce ZeroTuning, a training-free inference-time method that calibrates a single, task-agnostic token.
+Across a wide range of tasks and models, ZeroTuning improves performance over vanilla inference and is competitive with (and often stronger than) prior inference-time tuning baselines that require manipulating additional tokens or task-specific heuristics.
+ZeroTuning supports both supervised calibration (via a small labeled set) and unsupervised calibration (via entropy minimization), and it can be implemented across different optimized attention kernels.
+Overall, our results highlight the initial token as a simple but effective control point for lightweight, deployment-friendly model adaptation.
 Acknowledgments
@@
 Zhang et al.
-(2023a) introduces PASTA, which allows manual designation of important tokens during inference.
+(2023a) introduce PASTA, which allows manual designation of important tokens during inference.
 This is extended by AutoPASTA (Zhang et al., 2024a), which uses LLMs to autonomously identify salient tokens and increase attention to them.
@@
 PAI (Liu et al., 2024b) enhances attention to image tokens at inference time to counteract text-dominant bias.
-IBD (Zhu et al., 2024) and OPERA (Wei & Zhang, 2024) further refine this idea by prioritizing visual information or penalizing overconfident summary tokens.
+IBD (Zhu et al., 2024) and OPERA (Wei and Zhang, 2024) further refine this idea by prioritizing visual information or penalizing overconfident summary tokens.
 While effective, these methods depend on identifying task-specific tokens, which may introduce bias (e.g., overemphasizing misleading tokens) and limit applicability when token importance is unclear or attention maps are unavailable.
@@
 Xiao et al.
-(2023) shows that preserving such tokens is critical for maintaining performance in sliding window attention.
+(2023) show that preserving such tokens is critical for maintaining performance in sliding window attention.
 Kaul et al.
-(2024) attributes this effect to softmax normalization and causal masking, while Gu et al.
+(2024) attribute this effect to softmax normalization and causal masking, while Gu et al.
 (2024) and Barbero et al.
@@
 In this work, we provide a detailed analysis of attention tuning of the initial token across layers and heads, demonstrating its consistent influence across different tasks.
-Our approach bridges the gap between these lines of research by proposing a novel method that advances interpretable attention tuning.
+Our approach bridges the gap between the these lines of research by proposing a novel method that advances interpretable attention tuning.
 Appendix B Detailed Experimental Setup
@@
 Eager computes the full attention map; SDPA uses PyTorch’s efficient API to select the optimal implementation; Flash relies on fused CUDA kernels from the FlashAttention library.
-Tasks and Datasets: Our experiments encompass three task types across 15 datasets: (1) Text Classification and Reasoning, including SST-2 (binary sentiment classification) (Socher et al., 2013), SST-5 (fine-grained sentiment analysis) (Socher et al., 2013), MR (movie review polarity detection) (Pang & Lee, 2005), SUBJ (subjectivity classification) (Pang & Lee, 2004), TREC (question type classification) (Li & Roth, 2002), CB (commitment detection) (De Marneffe et al., 2019), and BoolQ (boolean question answering) (Clark et al., 2019); (2) Domain-Specific Multiple-Choice, including MMLU (cross-domain knowledge testing) (Hendrycks et al., 2020), AQUA (math word problems) (Zheng et al., 2024a), MathQA (algebraic reasoning) (Amini et al., 2019), LogiQA (logical reasoning) (Liu et al., 2023), CQA (commonsense reasoning) (Talmor et al., 2018), PIQA (physical commonsense QA) (Bisk et al., 2020), and ARCC (scientific reasoning) (Clark et al., 2018); and (3) Multi-Round Conversation, using MT-Bench (Zheng et al., 2024a).
+Tasks and Datasets: Our experiments encompass three task types across 15 datasets: (1) Text Classification and Reasoning, including SST-2 (binary sentiment classification) (Socher et al., 2013), SST-5 (fine-grained sentiment analysis) (Socher et al., 2013), MR (movie review polarity detection) (Pang and Lee, 2005), SUBJ (subjectivity classification) (Pang and Lee, 2004), TREC (question type classification) (Li and Roth, 2002), CB (commitment detection) (De Marneffe et al., 2019), and BoolQ (boolean question answering) (Clark et al., 2019); (2) Domain-Specific Multiple-Choice, including MMLU (cross-domain knowledge testing) (Hendrycks et al., 2020), AQUA (math word problems) (Zheng et al., 2024a), MathQA (algebraic reasoning) (Amini et al., 2019), LogiQA (logical reasoning) (Liu et al., 2023), CQA (commonsense reasoning) (Talmor et al., 2018), PIQA (physical commonsense QA) (Bisk et al., 2020), and ARCC (scientific reasoning) (Clark et al., 2018); and (3) Multi-Round Conversation, using MT-Bench (Zheng et al., 2024a).
 Baselines and Evaluation Metrics.
@@
 For an initially incorrect, uncertain sample, this often induces a ”competing peaks” phenomenon.
-Both Figure 12 and 13 illustrate cases where the model first amplifies an existing bias, leading to a deep entropy well corresponding to an incorrect answer.
+Both Figure 11 and 12 illustrate cases where the model first amplifies an existing bias, leading to a deep entropy well corresponding to an incorrect answer.
 Only with further increases in $\alpha$ does the model’s interpretation ”flip” to the correct one.
@@
 In these cases, the entropy minimum correctly corresponds to the right answer.
+Why Collective Signal Works Better?
+The visualizations in Figure 15 reveal a key insight: while individual uncertain samples exhibit highly irregular and sometimes multi-peaked probability/accuracy trajectories as $\gamma$ varies, the aggregate signal constructed from many samples becomes smooth, stable, and strongly aligned with true accuracy improvement.
+(a) Uncertain Population ($p<0.5$)
+(b) Certain Population ($p>0.5$)
+(c) All Samples (Aggregated)
+Figure 15: Population-level probability and accuracy trajectories. Gray lines represent individual sample trajectories. Red: average probability of the top token. Blue: average accuracy.
+We observe two complementary mechanisms:
+  • Statistical Cancellation for $\gamma<1$.
+Uncertain samples tend to fall into the format trap and display artificially low entropy for invalid outputs, whereas certain samples consistently lose confidence when $\gamma$ decreases.
+When averaged, these opposing trends cancel out, producing a flat and stable population-level curve.
+This prevents the collective entropy signal from being dominated by pathological samples.
+  • Signal Reinforcement for $\gamma>1$.
+In this region, both populations—despite their very different behaviors at the individual level—exhibit a shared upward trend in accuracy as $\gamma$ increases.
+This alignment creates a resonant effect: the aggregated probability and accuracy curves become smoother and strongly monotonic, reliably pointing toward the globally beneficial regime of $\gamma>1$.
+Together, these effects explain why Unsupervised ZeroTuning succeeds when optimizing the average entropy.
+The population average filters out idiosyncratic traps (format errors, bias amplifications) and amplifies the underlying structural signal that reflects the model’s latent reasoning and instruction-following capabilities.
+This makes the minimum of the average-entropy curve over $\gamma$ a trustworthy indicator of the optimal tuning direction.
 Appendix F LLM Output Example